### v1

# On the handling of method failure in comparison studies

Abstract
Comparison studies in methodological research are intended to compare methods in an evidence-based manner, offering guidance to data analysts to select a suitable method for their application. To provide trustworthy evidence, they must be carefully designed, implemented, and reported, especially given the many decisions made in planning and running. A common challenge in comparison studies is to handle the “failure” of one or more methods to produce a result for some (real or simulated) data sets, such that their performances cannot be measured in those instances. Despite an increasing emphasis on this topic in recent literature (focusing on non-convergence as a common manifestation), there is little guidance on proper handling and interpretation, and reporting of the chosen approach is often neglected. This paper aims to fill this gap and provides practical guidance for handling method failure in comparison studies. In particular, we show that the popular approaches of discarding data sets yielding failure (either for all or the failing methods only) and imputing are inappropriate in most cases. We also discuss how method failure in published comparison studies—in various contexts from classical statistics and predictive modeling—may manifest differently, but is often caused by a complex interplay of several aspects. Building on this, we provide recommendations derived from realistic considerations on suitable fallbacks when encountering method failure, hence avoiding the need for discarding data sets or imputation. Finally, we illustrate our recommendations and the dangers of inadequate handling of method failure through two illustrative comparison studies.
1 Introduction
In methodological research, comparison studies aim to compare the performance of methods to provide empirical evidence on their behavior and help data analysts choose appropriate methods for their setting. To be reliable, they require careful planning of the design, execution, and reporting of the results. These issues have gained increasing attention in recent years, both in the context of simulation studies based on artificially generated data (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024) and in the context of real data-based benchmarking studies (Boulesteix et al., , 2017; Friedrich and Friede, , 2024).
In this work, we focus on a particular aspect of the design of comparison studies that affects both simulation and benchmark studies, namely “method failure”. Method failure occurs when a method under investigation fails to produce an output for a (real or simulated) data set considered in the comparison study. The issue of method failure in comparison studies has begun to attract attention in recent years (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024), often focusing on non-convergence. However, there is little guidance on how to handle method failure appropriately, whether it occurs as non-convergence or in another form. This is illustrated by our review of published comparison studies from classical statistics and predictive modeling in Section 2 which shows that (reported) handlings often differ greatly between comparison studies. Nevertheless, these studies are positive examples of the reporting and the handling of method failure as this issue often goes unreported in many other comparison studies. For instance, in a systematic review of $42$ published simulation studies on the analysis of complex longitudinal patient-reported outcomes data, less than half “acknowledge that model non-convergence might occur” (Hinds et al., , 2018). Even more notably, only $12$ of $85$ applicable research articles reviewed by Morris et al., (2019) reported convergence as a performance measure. This contrasts with our perception—both through our own experiences and informal discussions with colleagues—that the problem of method failure is highly prevalent. In particular, signs of method failure are often traceable in the code provided for reproducibility even when not reported in the manuscript. This is often clear through the use of R functions try() or tryCatch() for the handling of errors, for example.
To illustrate the difficulty of handling method failure adequately, imagine conducting a comparison study in statistics or predictive modeling to compare the performance of a selection of methods. In the context of classical statistics, you likely focus on comparing the methods’ performances in “estimating one or more population quantities” (Morris et al., , 2019), sometimes called “estimand(s)” and base your analysis on data simulated using a certain data-generating mechanism (DGM). In predictive modeling, you are likely to compare the methods’ predictive performances, that is, their ability to make predictions on a new set of observations. Here, the analysis is typically performed on real benchmarking data sets. For simplicity, we refer to both settings as “simulation studies” and “benchmark studies”, respectively (though admitting that there may be intermediate situations, e.g. simulation studies investigating predictive performance on artificial data).
During your experiments, you encounter the failure of a method in one or more of the following ways: (i) you receive an error message or output of NA (“not applicable”) or NaN (“not a number”), implying that a certain calculation could not be completed (e.g., non-convergence), (ii) your system crashes, or (iii) the experiment runs for an excessively long time, forcing you to abort.
Table 1: Overview table of (a) a fictive simulation study and (b) a fictive benchmark study encountering method failure. This results in undefined estimates values for the simulation study (a) and undefined performance values for the benchmark study (b), marked in blue. In both cases, this complicates the aggregation of performance of the given method.
(a) Fictive simulation study: overview of estimates produced by three methods under investigation (columns) across $4$ repetitions, i.e. simulated data sets (rows), in one simulation scenario, where the true value is equal to $4$. An estimates value is obtained for each of the three methods and four simulation repetitions. Bias is commonly calculated as the mean deviation of the estimates from the true value across the $4$ repetitions. Method failure first leads to an undefined estimates value in the given repetition, which then complicates the estimation of performance of the affected method.
Repetition Method 1 Method 2 Method 3
1 3.89 4.23 4.08
2 3.78 4.13 4.11
3 NA 3.69 4.23
4 3.75 4.24 NA
Bias ? 0.07 ?
(b) Fictive benchmark study in predictive modeling: an estimated accuracy is obtained for each of the four data sets (rows) and three methods under investigation (columns). Method failure directly leads to an undefined performance (i.e., accuracy) value for the given method in the given row of the table, complicating the aggregation of its performance over all rows required to obtain an overall performance estimation.
Benchmark data set Method 1 Method 2 Method 3
1 0.85 0.88 0.87
2 0.9 0.91 0.86
3 NA 0.80 0.82
4 0.78 0.76 NA
Average accuracy ? 0.84 ?
Table 1 depicts a hypothetical summary of the results of your study in each simulation repetition/benchmark data set, denoting the undefined values by a blue “NA”. You may now wonder how to accurately determine the methods’ performances from this incomplete information. Note that Table 1 first of all highlights an important distinction that must be made between simulation and benchmark studies regarding the consequences of method failure. In a simulation study focusing on assessing the methods’ performances in estimating an estimand, a method typically produces a single estimate for the entire data set in the given repetition (i.e., row in Table 1(a)). Its performance is estimated as an aggregate over all rows in Table 1(a). In this setting, method failure first leads to an undefined estimate value which then complicates the assessment of the affected method’s performance. In a benchmark study focusing on the predictive performance of methods, on the other hand, a method makes a prediction for each row of the corresponding test data for a given benchmark data set (based on the decision rule generated on the training data). The method’s performance (e.g., accuracy) is then measured directly in each real data set (i.e., each row in Table 1(b)) and its overall performance is averaged over all rows of Table 1(b). Method failure results directly in an undefined performance value.
Looking at Table 1, it is tempting to view method failure as missing data: in each case, the row summarising results contains “holes” where results were expected. We caution against this. Missing values are data that exist and were intended to be recorded but for one reason or another are “not observed” (Rubin, , 2004; Carpenter et al., , 2023). In our context, although we intended to record a value, method failure means that the value does not exist for that particular data set. This is similar to data truncation by death as opposed to the inability to collect data due to the non-appearance of a patient at the doctor’s office (missing data). As a result, some of the most commonly used approaches to handling method failure such as imputation, being “missing data inspired”, are inadequate in most cases. Furthermore, only few are aware that the typical ways and reportings of discarding data sets associated with failure (either for all or the failing methods only) do not enable meaningful method comparisons. We illustrate this in detail in Sections 3 and 4.
In the subsequent part of this paper, we use the term “undefined” instead of “missing” to emphasize the important distinction between method failure and the “regular” missing data context. The consequence of our dissuasion from “missing data-inspired” approaches is that our work focuses intensively on alternative ways to deal with method failure.
The rest of this paper is structured as follows. In Section 2, we review and summarize the different manifestations of method failure in published comparison studies. In particular, we discuss how at the core, they often result from a complex interplay of multiple factors, and that an adequate handling requires thorough investigation on a case-by-case basis. Building on this, Section 3 is a compilation of practical recommendations for handling method failure which differentiates between different scopes of the study and research stages in which it may take place. In particular, we promote using fallback strategies when encountering method failure. This approach, if based on thorough and realistic considerations, enables the aggregation of performance across data sets even when method failure occurs. However, considering possible time constraints, we also discuss “slimmed-down” ways in which authors can still implement our recommendations. Finally, in Section 4, we demonstrate our recommendations using two fictive comparison studies that illustrate the complexity of the discussed issues.
2 Method failure in the literature
In our own experience, method failure occurs frequently when conducting comparison studies in statistics and predictive modeling. As already described, a look at the literature suggests that others also encounter this problem, either because method failure is explicitly mentioned in the text or because signs of method failure can be found in the code provided by the authors for reproducibility. Section 2.1 briefly outlines how method failure manifests and is handled in a selection of published studies from the classical statistics and predictive modeling literature. In Section 2.2, we argue that at the core, method failure is a complex interplay of multiple factors, deserving particular attention. Finally, Section 2.3 briefly categorizes different procedures for handling method failure used in these studies and provides a first outlook on the recommendations provided in Section 3.
2.1 Reporting and handling of method failure from the literature: vignettes
The following comparison studies should be viewed as positive examples, being among the few that report the occurrence of method failure, often alongside reasons and strategies for handling the resulting undefined values when summarizing and interpreting the results. The strategies adopted in these studies inspire—but do not completely overlap with—our recommendations given in Section 3. In these cases, we do not intend to criticize the applied approaches but aim to further improve the handling and reporting of method failure for future work.
van Smeden et al., (2016) investigate why different simulation studies offer conflicting minimal “events per variable” recommendations for logistic regression. They particularly observe that separated data sets (i.e., outcome variable is perfectly predicted by one or a combination of multiple covariates) are commonly omitted from the analysis since they lead to convergence issues of the likelihood maximization process. However, van Smeden et al., (2016) demonstrate that the removal of these data sets, in which the association between the covariates and outcomes is particularly strong, substantially affects bias, mean squared error (MSE), and confidence interval width, even if the frequency of separated data sets is small.
In Voeltz, (2024), a simulation study is conducted to compare three frequentist methods for the meta-analysis of diagnostic accuracy studies regarding sensitivity and specificity. Two of the three methods under investigation are occasionally affected by method failure due to non-convergence (while the third one does not employ an iterative process). In the case of non-convergence, one of the two does not return an error but outputs the current, non-converged estimate. As these outputs are often used in practice nevertheless, the authors also include them in the performance comparison. For the remaining method affected by non-convergence, they discard the data sets for which it fails and use the proportion of simulation repetitions in which the methods converge as an additional performance measure.
An extreme example of excluding performance data is to remove an entire method. This is done by Ruxton and Neuhäuser, (2013) in a simulation study to compare the empirical coverage of methods commonly used to calculate confidence intervals for the odds ratio. Given a binary outcome and binary exposure variable, some methods fail due to division by 00 if the corresponding $2\times 2$ contingency table contains one or more zero-entries.
Masaoud and Stryhn, (2010) compare the performance of different marginal and random effects estimation methods for the analysis of binary longitudinal data (e.g., GEE, Bayesian Markov Chain Monte Carlo) in a simulation study. With data simulated using two different data-generating mechanisms (i.e., stemming from marginal and random effects models), methods are compared regarding bias and coverage of the confidence intervals (as well as efficiency for data simulated from the marginal model). The authors encounter instances of “non-convergence or non-sensible estimates” (Masaoud and Stryhn, , 2010), for which the handling varies for the different methods affected. They either (i) change the software from R to SAS, (ii) switch the optimization technique used internally by the method, or (iii) for GEE, resort to using model-based standard errors if the method fails to produce meaningful robust standard errors. However, if method failure cannot be solved using these measures, they discard the corresponding data set.
Dunias et al., (2024) conduct a simulation study to compare the performance of four hyperparameter tuning procedures (including $k$-fold cross validation and the one-standard error rule for tuning) for prediction methods developed using predictive modeling methods such as Ridge, LASSO, or Random Forest. For some data sets, LASSO does not select any predictor variable, leading to a constant predictor. Resulting in the inability to assess the performance in terms of the calibration slope, the “[missing] calibration slopes [are] replaced by the maximum value of calibration slope for the corresponding scenario” (Dunias et al., , 2024). Note that this corresponds to imputation by the worst “observable” value since higher values of the calibration slope generally indicate that “risk predictions are too narrowly distributed” (Dunias et al., , 2024).
An approach to handling method failure by imputation is applied by Bischl et al., (2013) in their comparison study of $11$ methods for classification on $49$ real and simulated (benchmarking) data sets. They use misclassification error as performance measure combined with $100$ subsampling (also called “Monte Carlo cross-validation”) repetitions. When encountering method failure, they impute based on the proportion of repetitions in which the affected method fails. If below $20\%$, they impute by sampling from a normal distribution that is estimated from the remaining performance values of that model. This way of imputing serves to adhere to the assumptions of the model used to test the generalization performance. If a method fails for more than $20\%$ of repetitions, they impute by the worst possible value with the reasoning that the method’s behavior is “too unreliable for the current data set” (Bischl et al., , 2013).
Herrmann et al., (2021) perform a benchmark study to compare the predictive performance of survival time prediction methods based on high-dimensional multi-omics data using real cancer data sets. They encounter method failure in several cross-validation folds for several methods (including one method with excessive runtimes). Inspired by Bischl et al., (2013), they impute undefined performance values based on the proportion of “failed” cross-validation folds. If it is above $20\%$, they impute by the “worst possible value” of the corresponding metric, else, they impute by the mean performance of the method in the remaining “successful” folds. Note that the latter is equivalent to discarding those data sets in the performance assessment for which the given method fails. In a subsequent project, Nießl et al., (2022) investigate the effect of different imputation procedures on the performance assessment and ranking of the methods. They consider alternative procedures including imputation by the worst possible value and by a value that weights the method’s overall proportion of undefined performance values and its performance in the remaining folds. They show that the approach to imputing the undefined performance values has a notable effect on the performance ranking between the methods.
Another example can be found in the work of Hornung et al., (2024). They assess the performance of prediction methods that can handle block-wise multi-omics data by artificially inducing specific patterns of block-wise missing values into initially complete multi-omics data. Three of the considered methods fail for some data sets (with a maximum proportion of $34\%$). In some rare cases, failure is caused by memory issues. More commonly, however, the methods are “not applicable” (Hornung et al., , 2024) if the training and test data sets differ in the induced pattern of missing data blocks in certain ways. The resulting undefined performance values are addressed by discarding the corresponding data sets for all methods with the reasoning that discarding them only for the failing methods would lead to confounding by the induced pattern of block-wise missing values. Note that Hornung et al., (2024) argue that instances of memory issues are likely caused by a specific missingness pattern induced into the training data.
Fernández-Delgado et al., (2014) perform an extensive comparison study of supervised classification methods (“classifiers”) from several model families. The classifiers are assessed in $121$ (real) benchmarking data sets in combination with $4$-fold cross-validation and “accuracy” as performance measure (which is then averaged over all benchmarking data sets). Additionally, they compute the Friedman ranking (Sheskin, , 2011) based on accuracy. They encounter method failure in $2.1\%$ of classifier-data set combinations. As reasons for failure, they list “collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al., , 2014). Note that equal predictor values of all observations in the training data lead to the failure of the corresponding method since this violates its requirements. Memory issues also lead to method failure with large data sets, while particularly low numbers of observations in some classes in small data sets lead to errors for some classifiers. The authors discard the corresponding data sets for the failing methods to obtain each classifier’s accuracy. For the Friedman ranking, which requires an identical number of accuracy values for all classifiers, they impute undefined accuracy values with the mean value across all classifiers for this given data set.
Gijsbers et al., (2024) introduce a benchmarking tool for automated machine learning (“AutoML”) frameworks, which are considered as the “methods” under comparison. They conduct a comparison study of nine such methods based on various data sets using the Area Under the Curve (AUC) as performance metric. They categorize the reasons for failure into (i) exceeding available memory (or other memory-related issues), (ii) exceeding time limits, (iii) failure related to data set characteristics (such as highly imbalanced data), and (iv) failure caused by errors in the implementation of the AutoML framework. Categories (i) and (ii) are reported as the most common. Gijsbers et al., (2024) argue against ignoring undefined performance values, stating that they do not occur at random but are correlated with data set characteristics. They impute by the performance of a constant predictor.
2.2 Typical manifestations of method failure
Summarizing from the previous Section 2.1, there are common ways in which method failure often reveals itself to authors of comparison studies; see Table 2 for an additional overview.
Table 2: Common manifestations of method failure
Manifestation Explanation Examples from literature
Non-convergence
and other
calculation issues
Method cannot perform required calculations (within maximum number of iterations), therefore fails to produce a stable output. Cannot be solved by increasing memory or time budgets.
$\rightarrow$ Example: - division by zero - attempt to invert singular matrix $\rightarrow$ Can result in: - error message - non-meaningful output such as NA, NaN, output of $\infty$, or inflated estimation (for estimation task)
van Smeden et al., (2016)
Fernández-Delgado et al., (2014)
Memory
issues
Method consumes more memory than available to perform required computations.
$\rightarrow$ Can result in: system crashes (“fatal error”)
Hornung et al., (2024)
Fernández-Delgado et al., (2014)
Runtime
issues
Method fails to perform required computations within pre-specified or feasible time limits.
$\rightarrow$ “Failure by design” of study
$\rightarrow$ Real users would likely obtain valid output
Herrmann et al., (2021)
Gijsbers et al., (2024)
Non-convergence and other calculation issues  Authors often encounter the inability of a method to perform a calculation necessary for prediction or estimation, either implied through an error, a warning message, or a method output that is not meaningful. For methods employing an iterative procedure (e.g., an iterative maximum-likelihood procedure), this is often referred to as “non-convergence”, i.e., the method fails to produce a stable and valid output in the pre-specified number of iterations.
Memory issues  Even before exceeding possible time budgets, some methods crash due to memory issues. That is, the given method consumes more memory than is available. Gijsbers et al., (2024) additionally list segmentation faults as memory-related errors.
Runtime issues  A variant of method failure manifested and reported in existing comparison studies is excessive runtime. Note that in some comparison studies, fixed time budgets are predefined (possibly with an additional leniency period) such that exceeding them is declared method failure (Gijsbers et al., , 2024). Other studies do not predefine fixed budgets but find that, due to excessive runtimes, running the entire experiment for the affected method is impossible and must therefore be aborted (Herrmann et al., , 2021). Note that in contrast to the other two manifestations, method failure is actively declared by the authors when runtime issues occur. That is, running the methods under investigation on hundreds or thousands of data sets in the study creates the necessity of time restrictions, which are set by the authors arbitrarily, and without which the method would likely produce a valid output. In particular, in many cases, a corresponding real-world user, running the method on a single data set only, would likely obtain a valid output. This has important implications for the handling of method failure, which we will discuss in Section 3.2.
At first glance, it may seem like calculation, memory, and runtime issues are separate reasons for method failure, such that authors report specific “reasons” for failure for different methods within their study or even different “reasons” for a single method in different instances. For instance, some authors report non-convergence as a “reason” for failure in one method–data set combination and memory issues in another. However, upon closer examination, it becomes evident that these are not so much reasons but rather manifestations of method failure that are just the tip of the iceberg. In essence, method failure (or what might appear as such) is often the result of an individual and complex interplay of many different factors.
For example, Kriegel et al., (2017) show that runtimes can depend on numerous factors: “You might be surprised by the amount of differences in runtime that we point out, spanning up to several orders of magnitude in different implementations of one and the same algorithm. We found major performance differences even within the same toolkit, in different versions or releases: due to bugs, algorithmic improvement (having an indirect impact on the tested algorithm, e.g., in data handling), regressions, and environment changes. Finally, it is also noteworthy that good or bad choice of parameters for different algorithms can lead to void runtime comparisons”.
It is important to emphasize that differences in hardware (e.g., available memory) and/or software (e.g., operating system) can also affect assessment. As an extreme example of software dependency, consider that Herrmann et al., (2021) reported that one of the methods under investigation (sparse group lasso) was completely unusable due to an error in R that occurred using Windows as an operating system, while it did not occur under Ubuntu. As an extreme example of hardware dependency, consider that the feasibility of deep learning to train large artificial neural networks tremendously increases with the use of graphics processing units (GPUs). In essence, this means it is important to keep in mind that a computation that fails or is infeasible in terms of runtime and budget on one specific system setup, may work on a different setup. Note that the reverse situation that a computation that can be successfully conducted on one specific system cannot be conducted—i.e., reproduced—using a (slightly) different setup (e.g., during review), is very common and addressed by the term computational reproducibility.
In addition to these (combinations of) factors associated with a method (or its implementation) and the technical issues mentioned above, failure is typically also related to certain characteristics of the given data. That is, some characteristics can complicate required calculations for a method such that they are completely impossible or cannot be completed within the maximum number of iterations (non-convergence), or they exceed available time or memory budgets. For instance, van Smeden et al., (2016) report non-convergence of logistic regression algorithms when the outcome variable is perfectly separated by one or more predictor variables.
To sum up, oftentimes, certain characteristics of the data cause a certain method, in its current implementation, to fail in a certain computational environment.
2.3 Common approaches to handling method failure
The approaches to handling method failure from the published comparison studies in Section 2.1 can generally be divided into two categories. First, some authors resolve method failure so that no undefined values remain. This includes entirely removing failing methods from the study or modifying its implementation (e.g., adjusting some parameters) to make it work for all data sets. Second, other authors summarize the affected method(s) performance(s) despite undefined performance values. Most commonly, the corresponding data sets are discarded either for all or only for the failing methods. Other authors impute undefined performance values. Again, this is often done in various ways, for example, imputation through the worst possible value (when such a value exists) or through the worst (or mean) value achieved by other methods on the particular data set. Additionally, some authors also report the proportion of (real or simulated) data sets for which each method fails.
In the following Section 3, we discuss why imputation and discarding data (either for all or the failing methods only) are inadequate in most cases of method failure. In particular, we recommend alternative and more appropriate approaches that are based on realistic considerations regarding the research stage in which the study takes place and the corresponding targeted users.
3 Principles and recommendations for handling method failure
3.1 Absolute, relative, unconditional and conditional performance
There is no one-size-fits-all solution to handling method failure, exemplified by the examples in Section 2. This section discusses some principles to guide decisions. To illustrate ideas, we begin with two simple examples and some useful terminology. Suppose a simulation study aims to compare the performance of method A with method B in terms of bias. Method A fails in $20\%$ of repetitions while method B does not fail in any. Alternatively, suppose a benchmark study aims to compare the average accuracy of two methods. Again, method A fails in $20\%$ of benchmark data sets while method B does not fail in any. The hypothetical results for the two cases are given in Table 3(b) (a) and (b).
Table 3: Overview table of an illustrative comparison study comparing the performance of (a) two estimation methods regarding bias and (b) two prediction methods regarding average accuracy. Method A fails in $20\%$ of (a) simulation repetitions and (b) benchmark data sets.
(a) Hypothetical simulation study
Conditional
Bias Unconditional (repetitions where
(all repetitions) A does not fail)
Method A - 0.09
Method B 0.14 0.09
(b) Hypothetical benchmark study
Conditional
Average Unconditional (data sets where
accuracy (all data sets) A does not fail)
Method A - 0.85
Method B 0.72 0.85
We contrast the terms unconditional vs. conditional performance and absolute vs. relative performance. Unconditional performance simply means “in all data sets”, while conditional means “in data sets with no method failure”. The unconditional bias of method B is 0.14, while the unconditional bias of method A is not defined. The conditional bias of both methods is 0.09. Parallel descriptions apply to the hypothetical benchmark study in Table 3(b) (b). Absolute performance means “considered on its own” and relative performance means “when compared with another method”. In both Tables 3(b) (a) and 3(b) (b) the relative unconditional performance of methods cannot be quantified because the unconditional performance of method A is undefined. The relative conditional performance can be quantified: in Table 3(b) (a), the two methods have the same bias conditional on non-failure of method A; in Table 3(b) (b), the two methods have the same accuracy conditional on non-failure of method A.
These terms help to distill the problems with method failure in comparison studies. Firstly, the absolute conditional performance of method B should be of little interest. In deciding to use method B, why should we care about its performance when method A, a method we are not going to use, does not fail? We should care about a method’s unconditional performance. The problem is that we cannot compare the methods in terms of their unconditional performance. Comparisons should be performed vertically, or we are not comparing like with like: if we compared the unconditional performance of method B with the conditional performance of method A (which corresponds to discarding data sets for the failing methods), this would reflect poorly on method B, even though both perform equally well for the data sets without failure. This is all the more problematic since method B, in this case, is actually preferable to method A, as it outputs a result for all data sets!
The methods’ conditional performance (“discard corresponding data sets for all methods”) does compare like-with-like. In conducting this comparison, method B’s conditional performance is in both cases (a) and (b) better than its unconditional performance.
3.2 Caution with discarding data and imputation
It is clear that what we would really like to quantify, but cannot, is relative unconditional performance. The superficial solution often proposed when looking at results such as those in Table 3(b) is to consider undefined values due to method failure as missing values and use missing data methods, such as imputation to infer the unconditional performance of method A. After all, results given by different methods tend to be positively correlated, so results given by method B could inform the imputation.
We view it as mistaken to regard method failure as a form of missing data and have been careful to refer to “undefined” rather than “missing” values. The missing data literature defines missing values as those that exist but whose values were not recorded (Carpenter et al., , 2023; Rubin, , 2004). Method failure is far closer to what is called a truncating event in the estimands literature (Kurland et al., , 2009; Kahan et al., , 2024), which denotes a value that does not exist. That is, when a method fails, the result we wished to record simply does not exist for that data set. For this reason, we should be cautious about applying missing data approaches to method failure. In particular, imputing undefined performance values or discarding data sets associated with method failure (either for all methods or the failing methods only) do not enable the comparison of relative unconditional performance. There are two exceptions to this.
First is when a “failure” is recorded but a result could have been obtained; for example, if failure occurs due to excessive runtime but the definition of excessive in the comparison study can safely be judged as shorter than in practice. For cases like this, it may make sense to run the experiment on a randomly selected subset of all simulation repetitions/benchmarking data sets for the method affected by runtime issues (“failure by design”). As this can be seen as missingness mechanism “missing completely at random” (missingness is “unrelated to the observed and unobserved data” (Carpenter et al., , 2023)), discarding the data sets for which the method “fails” or imputing based on the performance results from other methods or data sets may be appropriate. To ensure missingness completely at random, it is important to select the subset of data independently of the data sets for which the method has particularly long runtimes. Researchers should therefore think about possible runtime problems and carry out the corresponding random selection of data sets before conducting the study (see Section 3.3.7).
A second exception is when failure occurs for no reason, such that it is safe to compare the unconditional performance of method B with the conditional performance of method A. However, because method failure always happens for a reason – the interplay of a particular method with a particular data set – we do not believe this would ever occur.
An often more suitable approach to handling method failure than discarding data sets or imputing, which we term the “fallback strategy”, involves specifying a backup method when a failure occurs. This strategy is described in more detail in Section 3.3.5.
Table 4: Recommendations on handling method failure
Caution with Explanation
Imputation In most cases: method failure does not lead to “missing data”
$\rightarrow$ Possible exception: Runtime issues
Discarding data
(for all methods or the
failing methods only)
Usually does not provide assessment and comparison of relative unconditional performance
$\rightarrow$ Possible exception: Runtime issues
Alternative recommendations Explanation
Pinpoint interplay causing failure as the basis of any of the following considerations…
Check for method misuse Is the failing method truly developed for given estimation or prediction task?
Check if data matches scope Are targeted users likely to encounter this kind of data and failure?
Be realistic when modifying method implementations Consider changes in implementations to solve method failure that are realistic for targeted users and research stage. Possibly combine with fallback strategies (see below).
Incorporate fallback strategy Incorporate method pipelines “method A if it succeeds in producing an output, method B otherwise”.
$\rightarrow$ Enables assessment of relative unconditional performance.
Report transparently … - failure proportions, ideally as performance measure. - interplay causing failure with sufficient detail: “Method Z, with an overall failure proportion of $X\%$, fails when … and when … .” - chosen handling(s) and justify choices based on above considerations.
Slimmed-down recommendations
Analyze and report simultaneously:
(i) results after discarding data sets for all methods.
(ii) results after discarding data sets for the failing methods.
(iii) proportions of method failure for all methods.
$\rightarrow$ Report that this does not provide assessment of relative unconditional performance.
3.3 Practical recommendations in the presence of method failure
In the following, we present practical recommendations in the presence of method failure that may be more appropriate than imputing or discarding data sets. We advise implementing these recommendations as comprehensively as possible as they contribute to increased usability of comparison studies to users of the methods. However, we are aware that methodological researchers facing time constraints and page limits may be prevented from following them completely. In Section 3.4, we therefore discuss how a “slimmed-down” version of our recommendations can be applied nevertheless.
3.3.1 Pinpoint interplay of method and data characteristic causing failure
We recommend authors of comparison studies investigate and understand for which types of data sets and why a corresponding method fails, beyond the initially perceived symptom(s). Ideally, they should identify if particular (and if different) data characteristics cause the failures, leading to evaluations such as “method A fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes”. Having identified the dynamic between the given method and data that causes method failure, authors should base their further handling on the considerations and recommendations provided in the following Sections 3.3.23.3.7.
3.3.2 Check for method misuse
When encountering method failure, authors should first check whether the method was truly developed for the given estimation or prediction task, and the type of data (based on the investigations from the previous Section 3.3.1), and if it is used as intended by the developers. If not the case, what appears initially as method failure may actually be method misuse. Method misuse should not be attributed to the method but to the user, that is, the method should not be penalized for failing on a data set or a task it was not intended for. To avoid method misuse and any associated failure in the first place, authors of comparison studies should familiarize themselves thoroughly with the methods under investigation and read the corresponding user manuals carefully. It may also make sense to consult the developers of the individual methods (and possibly involve them in the study).
Note that there may also be cases where a method is commonly misused by real-world users, especially when they are inexperienced. Depending on which kind of user the study aims at (e.g., inexperienced versus experienced), the occurrence of this kind of failure may be relevant to them and should therefore be included in the study. In these cases, the misuse should be clearly emphasized.
3.3.3 Check if data matches scope of study
If method failure remains after solving or ruling out method misuse, authors should assess if real-world users are likely to encounter such data (and therefore, this method failure) in practice. These considerations must include the scope of the study and are of course not purely objective. Importantly, they should be reported in the study, since they carry valuable information and affect the interpretation of the study’s results.
Consider the case of method failure caused by the method’s inability to make required calculations due to an “extreme” structure of the given simulated data set, for example highly imbalanced classes in the outcome or predictor variable(s), as in White et al., (2024). Depending on the scope of the study, these extreme data sets may be of interest or not. If they are deemed unrealistic for the considered scope and thus not relevant to the readers of the study, authors may consider making sensible parameter changes in the DGM to generate data that is more realistic given the aim of their study, in line with White et al., (2024). Similarly, “too extreme” data may also affect real data based studies. If one of the (real) benchmarking data sets is considered to be irrelevant to the scope of the study after careful examination, authors may consider removing the data set from the study completely. See Section 4.2 for an illustration.
However, if the data sets causing failure are deemed relevant to the study’s scope, the researchers conducting the comparison study should—more appropriate than discarding or imputing in most cases—proceed by considering how real-world users are likely to behave in this situation. In particular, they may modify certain parameters of the concerned method(s) (see Section 3.3.4) or resort to a fallback (see Section 3.3.5).
3.3.4 Be realistic when modifying method implementations
As already touched upon, in some instances of method failure, a way of mimicking the behavior of real-world users may be to resort to an alternative implementation of the affected method (see Masaoud and Stryhn, (2010) for an example). This way, method failure may be resolved completely, allowing for the assessment of unconditional (absolute and relative) performance of the methods under investigation.
For instance, one could modify a parameter of the method, such as increasing the maximum number of iterations in the iterative maximum likelihood process or changing the internally employed optimization algorithm. However, authors of comparison studies should not tweak the implementation of the method thoughtlessly. Instead, they should always consider if the applied change in the parameter setting fits the given research stage and corresponding scope of the study. For instance, if the study goal is to understand the methods’ behaviors quite generally (and in earlier phases of research according to the terminology of Heinze et al., (2024)), authors may modify a greater variety of parameters to see when the method is successful and when it fails. However, if the study takes place in later phases and aims to guide inexperienced users, the applied parameter changes should be more restricted to what a naive user of the method would be able and likely to do. In any case, the performance assessment of the given method is only meaningful if the authors stress and justify their parameter choices when reporting the study.
Furthermore, it is important to make sure that the method comparison is still fair, meaning that the effort spent on each method under consideration should not differ significantly between the methods, especially in late phase studies, where neutrality is of crucial importance. Finally, one should be careful about increasing the runtimes of the affected method(s) when running a comparison study. Especially when the number of iterations is raised, runtime or memory consumption can increase notably and might result in runtime or memory issues so that method failure is, after all, not solved.
If, after these considerations, a change in the implementation of the given method is deemed reasonable, any deviation from the default parameter setting of a given method should be reported and justified transparently. Note that, as will be discussed in Section 3.3.5, changes in a method’s implementation could also be applied as a fallback strategy.
3.3.5 Consider a fallback strategy
When contemplating ways to mimic real-world users when encountering method failure in later research stages, Morris et al., (2019) argue that they would usually not give up but naturally resort to an alternative method, a fallback, that successfully provides an output. Therefore, it can make sense to transfer this fallback strategy into the context of methodological comparison studies and evaluate the performance of method pipelines of the type
“use method A if it successfully produces an output, use method B otherwise”.
On the one hand, this approach incorporates what data analysts really do in practice, and on the other, it indirectly yields automatic handling of undefined performance values in case they affect method A but not method B. If chosen carefully and in a way that there is a suitable fallback for each instance of method failure, fallback strategies allow for the assessment of (absolute and relative) unconditional performance of the methods under investigation. Besides realistically reflecting the behavior of real-world users, this approach results in a performance assessment that depends only on the definition of the pipeline. Note that when using fallback strategies, we no longer evaluate the “pure” performance of method A. However, it is clear from Section 3.1 that this in an unachievable goal in the first place.
The concept of fallback strategies is already intuitively considered in some existing comparison studies. Voeltz, (2024), for instance, conclude from their study that “the approach by Frömke et al., (2022) could be used as a fallback strategy in case of non-convergence [of the overall best-performing method] because it always yields results”, where the approach by Frömke et al., (2022) is a non-iterative procedure, therefore never affected by non-convergence by design. Going one step further than Voeltz, (2024), one could provide a performance assessment for the resulting pipeline.
A special case of the fallback strategy has indirectly been addressed in Section 3.3.4: it may sometimes make sense to define an alternative implementation of the given method (e.g, using a different optimizer or an increased number of iterations) as a fallback instead of a completely different method (as in Crowther et al., (2014) who changed the number of quadrature points). The corresponding pipeline would then consist of the default implementation of that method for those data sets where it successfully produces an output and the alternative implementation for the remaining data sets.
In practice, the choice of a suitable fallback depends on the research phase. For instance, in early phases, the fallback may be constituted of an alternative implementation (e.g., alternative parameter choice) of the original method. Also, authors should strongly base the selection of fallbacks on what can be considered realistic given the expertise in statistics and programming of the targeted users. For instance, in a study targeting data analysts with limited statistical background, the fallback strategy may be an alternative method with default parameters rather than the same method with complex parameter settings. Since the choice of suitable fallbacks involves some subjectivity, authors should justify their choice(s) transparently.
If failures are expected at the planning stage, i.e., before conducting the study, one may consider the whole pipeline as a competing method to be assessed in the comparison study. Another approach would be to design a pipeline after seeing the results of the comparison study, as did by Voeltz, (2024). In this case, the definition of the pipeline can be viewed as a result of the comparison study rather than a method that has to be evaluated in the first place. The definition of the fallback strategy, however, may not always be as straightforward as in Voeltz, (2024). It may for example make sense to consider complex rules involving a second fallback in case the first one fails.
3.3.6 Transparently report failure proportions, underlying interplays, and applied approaches
Following all recommendations above, authors should transparently report (i) any occurrences and proportions of method failure, (ii) the underlying interplay(s) causing failure (in different instances), and (iii) the corresponding approaches following the in-depth investigations of method failure.
Report any occurrences and proportions of method failure  This can serve as an important criterion for real-world users selecting a suitable method. Even the “best performing” method is of limited use if the amount of settings in which it can be applied without frequent failures is limited. Ideally, the proportion of method failure should be included as an additional performance measure for all considered methods (see Morris et al., (2019)); see Voeltz, (2024) for an example. If method failure is predictable for certain settings and all affected methods fail for the same data sets (see, e.g., the example from Section 4.1), it may make sense to include the binary performance measure of the type “can handle setting: yes. vs. no” and report it alongside the corresponding number of repetitions/benchmark data sets associated with the failure.
Report interplay of method and data causing failure  Adding to the previous paragraph, a corresponding reporting for each method can look like “method A, with a failure proportion of $X\%$, fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes”. This enables a more comprehensive overview of the applicability of the methods compared to reporting the corresponding failure proportions only. Readers can learn in which settings a method is applicable, and no less importantly, in which it is not. Additionally, these insights serve as justification of the chosen handlings to method failure. An example of a corresponding reporting can be found in Hornung et al., (2024) who provide detailed descriptions in the supplement of their paper.
Report applied handlings  Any actions resulting from the above-described in-depth investigations and recommendations regarding method failure should be reported with sufficient detail. This includes the removal of any misused methods, the removal of data (or modification of DGMs) that do not match the scope of the study, any changes in implementations, and the incorporations of fallback strategies. Additionally, authors should justify their choices given the scope of the study (i.e., which users it is aimed at), and the research stage in which it takes place.
3.3.7 Finally: think ahead, pre-register but remain flexible
The previous examples and those that will be outlined in Section 4 make clear that method failure and its cause(s), related to complex dynamics between factors such as method and data set, are often difficult to anticipate. While pre-specifying the handling of undefined values has the advantage of reducing analytical uncertainty and its dangers in terms of selective reporting (Siepe et al., , 2024), pre-specifying a standard procedure prior to conducting the study may sometimes be delicate.
Our recommendation is, as far as possible, to consider in advance which combinations of data and methods could lead to failure. For instance, carefully reading user manuals and help pages of all methods before conducting the study might reveal specific requirements of a method on the data. This may make some instances of method failure less unexpected. More generally, making method failures less unexpected can be achieved by conducting a pilot experiment. Pilot experiments test whether an experimental design is appropriate for the specific context in which it is applied, and are recommended in several textbooks on Design and Analysis of Experiments (Cohen, , 1995; Dean et al., , 2017; Montgomery, , 2020). For example, in the context of real data comparison studies, a pilot experiment might use a fraction of the available data sets, say $20\%$, to test whether the specified method implementations produce reasonable results. The remaining $80\%$ of the data sets are then used for the main experiment, which may be refined based on what is learned in the pilot. A pilot study can also help authors anticipate possible runtime issues for which they can then pre-specify the random subset of data sets on which the experiment for the affected method will be restricted. Of course, it is important to clearly report the results of the pilot study too. After implementing these strategies, pre-specifying a procedure for handling method failure may be less hazardous. However, even despite pre-registration, researchers should be aware that unforeseen issues may appear and require adaptations of the pre-specified strategy. See Lakens, (2024) for a discussion of the consequences of deviations from pre-registration on the severity and validity of inferences.
3.4 Slimmed-down recommendations when time is short
For some authors, there may be time constraints that do not allow for any (in-depth) implementations of our recommendations from the previous Section 3.3, however, authors nevertheless wish to aggregate results over data sets despite the existence of undefined values caused by method failure. In these cases, where it is difficult to follow the entirety of our recommendations, we recommend the following slimmed-down analysis and reporting strategy. Authors may present performance results after (i) discarding the corresponding data sets for the failing methods, (ii) discarding these data sets for all methods, and (iii) report the failure proportions for all methods, ideally as an additional performance measure. This three-fold analysis and reporting enables more differentiated assessments of the methods’ performances and avoids the preference of methods based solely on how data sets associated with method failure are discarded. In particular, it provides insights into how the methods successfully producing an output perform on the data sets for which others fail (and how this can influence the method comparison).
However, following Section 3.1, authors should be aware and clearly state in their manuscript that any form of aggregating performance results despite existing undefined values does not allow an assessment of unconditional performance of all methods under investigation.
4 Illustrations
In the following, we illustrate the guidelines presented in Section 3 through two illustrative comparison studies featuring method failure: a statistical simulation study and a benchmark study at the interface between statistics and predictive modeling. For both examples, we first demonstrate how the ad hoc approaches discussed in Section 3.2, such as discarding data sets associated with failure, can distort the results and even lead to misleading method recommendations. Then, we illustrate the implementation of the recommendations from Section 3.3. Thereby, our focus is exclusively on the handling of undefined performance values and its consequences on the method comparison, rather than the method comparison itself. This has three consequences. First, both illustrations represent only a small excerpt (e.g., a few or a single scenario) of what would be a typical publishable study. Second, we do not give a comprehensive overview of all aspects of the studies but focus on those that are relevant to the issue of undefined performance values. Third, our goal is not to provide insights into the performance of the compared methods (larger dedicated studies would be necessary for this purpose).
4.1 Example 1: Comparing methods for the estimation of odds ratios
4.1.1 General
Our first example, a statistical simulation study, deals with the comparison of odds ratio estimation methods. Odds ratios (“ORs”) quantify the association between an exposure and an outcome and are typically used in (retrospective) case-control studies. Here, we focus on the case of a binary exposure $X\in\left\{0,1\right\}$ and binary outcome $Y\in\left\{0,1\right\}$. While a straightforward way of estimating the OR is to calculate $\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$ (where $n_{ij}$ is the observed number of subjects with $X=i$ and $Y=j$), other methods employ more complex modeling strategies. Note that ORs are bounded in $(0,\infty)$, where values $>1$ imply a positive association between exposure and outcome.
With the true OR value being generally unknown in real data sets, OR estimation methods are typically compared using simulation studies under different scenarios, e.g., different sample sizes, underlying true OR values, and probabilities of exposure $p_{x}$. Especially when the sample size is low, the underlying true OR is far from $1$, or $p_{x}$ is close to 00 or $1$, some simulated data sets may not contain any subjects for one or more of the four outcome-exposure combinations. This results in $n_{ij}=0$ for the corresponding combination $i,j\in\{0,1\}$ for this data set. These zeros are called “sampling zeros” and cause computational issues (e.g., division by 00) for many of the available OR estimation methods. They lead to either returning an error or, depending on the position of the sampling zero in the contingency table, non-meaningful OR estimations of 00 or $\infty$”, resulting in undefined performance values in the corresponding data sets. In practice, users often apply the “Haldane-Anscombe correction” (Haldane, , 1956; Anscombe, , 1956) to handle sampling zeros in the OR estimation. This correction consists of adding an offset of $+0.5$ to all four cells in the contingency table to avoid issues in computation.
4.1.2 Design of the fictive study
In our fictive simulation study, we vary the underlying true OR (OR $\in\{2,3,4,5\}$) and exposure probability $p_{x}$ ($p_{x}\in\{0.25,0.5\})$ while keeping the sample size at the fixed value of $n.obs=50$. This results in eight distinct simulation scenarios (see Table 5 for an overview). For each scenario, we simulate 100 000100000100\,000100 000 outcome-exposure data sets (where the exposure variable $X$ is generated using a Bernoulli distribution with probability $P(X=1)=p_{x}$ and outcome $Y$ is generated using the logistic function with the linear effect of exposure $X$ specified as $\beta_{X}=\log(\text{OR})$). Five OR estimation methods are applied to the simulated data to obtain a point estimate for the true OR, namely “Fisher”, “Midp”, “Small” (all three implemented in R package epitools (Aragon et al., , 2020)), “Woolf” (Woolf et al., , 1955) (implemented in R package pairwiseCI (Schaarschmidt and Gerhard, , 2019)), and “Manual”, which corresponds to the manual computation $\widehat{OR}=\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$. Note that we omit method “Wald” from the epitools package since its estimation output is identical to the manual computation (i.e., method “Manual”). From the five estimators under investigation, only Small and Woolf can handle sampling zeros internally. For each OR estimator and simulation scenario, bias on the $\log$-scale is computed as
[…]
and the MSE on the $\log$-scale is obtained as
[…]
Note that the $\log$-transformation ensures symmetry around the value of “no association”. For both performance measures, values closer to 00 indicate better performance (Morris et al., , 2019).
4.1.3 Occurrence and handling of method failure
Table 5 displays the proportion of the 100 000100000100\,000100 000 simulated data sets containing at least one sampling zero across the eight simulation scenarios. As expected, this proportion increases with an increasing true OR$>1$, while decreasing when exposure $X$ is more balanced.
Table 5: Overview of the proportion of the 100 000100000100\,000100 000 simulated data sets with sampling zeros across the simulation scenarios. $n.obs$: number of observations, $p_{x}$: exposure probability.
Simulation scenario n.obs true OR $p_{x}$ Sampling zero proportion
1 50 2 $0.25$ 1.28%
2 50 3 $0.25$ 3.96%
3 50 4 $0.25$ 7.8%
4 50 5 $0.25$ 12.0%
5 50 2 $0.5$ 0.01%
6 50 3 $0.5$ 0.11%
7 50 4 $0.5$ 0.50%
8 50 5 $0.5$ 1.27%
Our analysis consists of two parts. First, we compare the performance of the OR estimation methods regarding performance measures bias and MSE when ignoring data sets with sampling zeros for all methods versus ignoring data sets only for the failing methods (“ad hoc approaches”). While we are aware that users often apply the Haldane-Anscombe correction when encountering sampling zeros, this first part illustrates how the preference of some methods may mainly be driven by the way data sets with sampling zeros are discarded.
Second, we repeat the analysis implementing the principle of fallback strategies suggested in Section 3.3.5. As fallbacks for methods Midp and Fisher for data sets with sampling zeros, we (i) either apply the Haldane-Anscombe correction, (ii) use method Small, or (iii) use method Woolf. For the manual estimation of the OR (method “Manual”), we only apply the Haldane-Anscombe correction as a fallback, assuming that a user applying manual computation in the first place, which is very easy to implement, is unlikely to resort to a method that requires notably more effort. Together with Woolf and Small, nine “methods” (“pure” methods or pipelines) are compared in total and the results are presented in Section 4.1.4.
Note that since we focus on the effect of handling method failure (instead of the actual results of the comparison), and this effect becomes clearer through examining the individual scenarios, we do not aggregate the performances across the eight scenarios in either of the two analysis parts.
Figure 1: Performance ranks of the OR estimation methods across the eight simulation scenarios, derived from the empirical bias in subfigure (A) and from the MSE in subfigure (B). In both subfigures, the ranks are compared between discarding data sets yielding failure for the failing methods only (denoted by “Failing” on the $x$-axis) versus for all methods (denoted by “All” on the $x$-axis). A rank of 1 means that the method has the lowest empirical bias (A) or MSE (B) in the given scenario. Note that Small and Woolf are the only methods in this selection that can deal with sampling zeros internally. Abbreviation of methods: “Fis”: Fisher; “Mid”: Midp“; “Sma”: Small; “Woo”: Woolf; “Man”: Manual (i.e., manual estimation of the OR).
Ad hoc approaches  For the first analysis, Figure 1 shows that the way of discarding data sets with sampling zeros can substantially influence the results of the method comparison. This particularly affects estimator Woolf, one of the two methods that can handle sampling zeros internally. In terms of bias (Figure 1 (A)), its performance rank deteriorates in most scenarios and by up to three positions when discarding data sets with sampling zeros for all methods (including Woolf and Small). The reason for this is that across the simulated data sets without sampling zeros, Woolf slightly underestimates the true OR, while for the few data sets with sampling zeros, Woolf notably overestimates it. This leads to top performance when including all data sets in its performance assessment (while discarding data sets with sampling zeros for the failing methods). Even more so, the manner of discarding data sets with sampling zeros affects Woolf’s performance rank regarding MSE (Figure 1 (B)). In some scenarios, Woolf’s rank improves from bottom to top rank, indicating its poorer performance for data sets with sampling zeros. The performance rank of Small, the second method that can handle sampling zeros internally, is less affected by this for both bias and MSE. In particular, it scores last in terms of bias across all scenarios and for both manners of discarding (Figure 1 (A)). However, its performance rank changes by up to two positions for MSE (Figure 1 (B)). The performance ranks of the remaining methods are also affected regarding bias and MSE, where the manner of discarding data sets with sampling zeros affects their performance ranks by up to two positions.
Note that the manner of discarding data sets has a larger impact on the performance comparison for scenarios with a higher proportion of sampling zeros. In particular, scenarios 5 and 6, with a maximum of $0.11\%$ of data sets containing sampling zeros, yield equal performance rankings across both handlings. However, differences in rankings within a given scenario are already observable with a proportion of $0.5\%$ of data sets with sampling zeros.
Figure 2: Performance ranks of the OR estimation methods across the eight simulation scenarios when applying fallback strategies to methods Midp (“Mid”), Fisher (“Fis”), and Manual (“Man”). The rankings are derived from the empirical bias in subfigure (A) and from the MSE in subfigure (B). A rank of 1 means that the method has the lowest empirical bias (A) or MSE (B) in the given scenario. Note that for Midp and Small, the fallbacks Small (“Sma”), Woolf (“Woo”), and the Haldane-Anscombe correction (“$+0.5$”) are used. For Manual, only the Haldane-Anscombe correction is used as fallback. Small and Woolf are the only methods in this selection that can deal with sampling zeros internally, therefore not requiring any fallback. This amounts to nine “methods” in total. The combination of original method $m_{1}$ and fallback $m_{2}$ is marked by “$m_{1}/m_{2}$”.
4.1.4 Recommended handling
As expected, the relative performance between Woolf and Small remains the same between applying fallbacks for the other methods (Figure 2) versus discarding data sets with sampling zeros for the failing methods only (“failing” parts in Figure 1 (A) and (B)). This makes sense since neither method requires a fallback, resulting in unchanged bias and MSE. Midp’s overall good performance regarding bias and MSE is underpinned by using fallbacks Small and the Haldane-Anscombe correction. Here, it is debatable which fallback is preferable. Both fallbacks for Midp frequently switch ranking positions across the performance measures and simulation scenarios. However, fallback Small might be even easier to implement than the Haldane-Anscombe correction, requiring only a single parameter modification from Midp. Note the interesting observation that Small scores below the remaining methods regarding bias when applied alone, but performs well as a fallback for Midp.
4.2 Example 2: Comparing methods for computing confidence intervals for the generalization error
4.2.1 General
The fictive study aims to compare the performance of two methods for constructing confidence intervals (CIs) for the generalization performance of a prediction model based on resampling-based estimates, such as cross-validation. The “naive” method “$N$” for constructing the CIs ignores the dependence between the generalization performances in the resampling repetitions, while method “$C$” corrects for it. The performance of both methods is assessed in terms of coverage, which is the probability that a constructed CI includes the true value of the generalization performance. It should be close to its nominal value, e.g. $0.95$ if we consider $95\%$-CIs.
4.2.2 Design of the fictive study
Classification trees (Breiman et al., , 1984) are considered as models for predictive modeling. As benchmark data set, we use the data set provided by the SAPA project (Condon et al., , 2017), further processed according to Klau et al., (2023). The response variable is binary and five variables from the data set are used as predictors. The generalization performance is quantified using the Area Under the Curve (AUC), where a value of $0.5$ corresponds to a random prediction and a value of $1$ indicates perfect classification. The AUC is estimated using $15$ repetitions of repeated subsampling ($i=1,\ldots,15$) with a 4:1 ratio (i.e., the data is randomly split $15$ times into $80\%$ on which the model is trained and $20\%$ on which the AUC is obtained (Hastie et al., , 2009)).
Method $C$ suggested by Nadeau and Bengio, (2003) computes the CI for the generalization AUC as
$\Big{[}\ \overline{\widehat{AUC}}-t_{0.975,14}\cdot\sqrt{\Big{(}\frac{1}{15}+c\Big{)}\cdot S^{2}_{\widehat{AUC}_{i}}},\quad\overline{\widehat{AUC}}+t_{0.975,14}\cdot\sqrt{\Big{(}\frac{1}{15}+c\Big{)}\cdot S^{2}_{\widehat{AUC}_{i}}}\ \Big{]},$ (1)
where $\overline{\widehat{AUC}}$ represents the mean AUC estimate over the $15$ subsampling repetitions and $S^{2}_{\widehat{AUC}_{i}}$ the corresponding sample variance, while $t_{0.975,14}$ is the $95\%$-quantile of the $t$-distribution with $15-1=14$ degrees of freedom. The fixed correction term $c>0$ incorporates the correlation structure between the subsampling repetitions (Nadeau and Bengio, , 2003). Method $N$, on the other hand, does not apply any correction, i.e. sets $c=0$ in the above formula.
[…]
The study is conducted using R package mlr3, method $N$ is implemented using function t.test() from package base and for method $C$, $\overline{\widehat{AUC}}$ and $S^{2}_{\widehat{AUC}_{i}}$ are obtained using functions mean() and sd() (from the same package).
Figure 3: Design of the fictive study to obtain the empirical coverage of methods $C$ and $N$. An approximation of the “true AUC” is obtained based on the very large $D_{test}$.
4.2.3 Occurrence and handling of method failure
For method $N$, the implementation of the above-described procedure returns the following error message for 300 of the $1000$ iterations in R:
Error in t.test.default: data are essentially constant.
Hence, the confidence intervals cannot be constructed for method $N$ in these iterations, making it impossible to assess whether they cover the true AUC. Method $C$ is not affected by these errors. At first glance, the error message gives little indication of how to resolve it. In the following, we first apply three ad hoc approaches to handle these errors. Then, by following our in-depth guidelines from Section 3, we demonstrate how these ad hoc approaches prove to be inadequate, and illustrate more suitable ways to handle method failure in this example.
Ad hoc approaches  One could be inclined to apply one of the following ad hoc approaches:
  • 1.
    for method $N$, ignore all data sets in which it fails (method $C$ is unaffected).
  • 2.
    for both methods, ignore those iterations in which $N$ fails.
  • 3.
    for method $N$, set “undefined” confidence intervals to “not covering the true AUC”.
Note that the intuition to approach 3, which can be viewed as a form of imputation, is that a CI that does not exist cannot cover any value. It thus induces a negative correlation between the proportion of undefined confidence intervals and the coverage of method $N$.
Table 6: Empirical coverages of methods $N$ and $C$ resulting for (a) the three “ad hoc” approaches and (b) setting undefined CIs to zero-width CIs.
(a) Ad hoc approaches: NA-CIs are (b) NA-CIs are
discarded for
method $N$ only
discarded for
methods $N$ and $C$
set to “not covering
true AUC”
set to zero-width
interval
Method $N$ $0.54$ $0.54$ $0.38$ $0.68$
Method $C$ $1$ $1$ $1$ $1$
The coverages resulting from the ad hoc approaches are displayed in Table 6 (a). While the coverage of method $C$ is constant across all three ad hoc approaches (coverages above the nominal level indicate conservativeness), method $N$, while too liberal across all approaches, is even more liberal using the ad hoc imputation approach.
4.2.4 Recommended handling
We now closely inspect the modeling processes underlying the study. The goal is to find out the root cause of the error message and to adapt the handling of the undefined performance values accordingly. The errors in method $N$ occur when the estimated AUCs are equal across all $15$ subsampling repetitions, resulting in a sample variance $S^{2}_{\widehat{AUC}{i}}$ of 0. This zero variance in the denominator of the $t$-statistic causes an error in function t.test(). However, focusing on the confidence interval, we realize that from a theoretical point of view, a sample variance of 0 simply means that the CI is the zero-width interval $\left\{\overline{\widehat{AUC}}\right\}$. Therefore, method $N$ not yielding a CI (and returning an error) is a technical artifact of its implementation. A straightforward solution would be to modify the implementation of method $N$, using mean() and sd() instead of t.test(), so that it computes the zero-width CI, which can enter the calculation of coverage as any other CI (see Table 6 (b) for the resulting coverage).
However, taking the investigation another step further reveals that a sample variance $S^{2}_{\widehat{AUC}{i}}=0$ always occurs alongside all $15$ estimated (and the approximate true) AUC being equal to $0.5$. Ultimately, this worst possible performance of the classification tree (“random prediction”) occurs when the classification tree does not perform any splitting. This is related to the characteristics of the data set. For example, the problem disappears if we include additional predictors: splitting then occurs, leading to an AUC higher than $0.5$.
The adequate handling of method failure in this illustrative study, as described in Section 3, depends on the aim of the study. In particular, a prediction problem in which splitting often does not take place can be viewed as an extreme scenario. If the study’s scope includes such extreme scenarios, one could attempt to suitably modify parameters of the classification tree to favor splitting. If, however, the study aims to compare both methods in regular (i.e., less extreme) scenarios, one might consider additional predictors to favor splitting more strongly, resort to an alternative prediction model (e.g., logistic regression), an alternative measure of generalization performance (e.g., accuracy), or use a different benchmarking data set.
Finally, note that this example illustrates that method failure may be complex and that the interplay causing it may not be immediately identifiable, even if the benchmark study consists of a single scenario only. In a real benchmark study which typically contains multiple scenarios, thorough investigations can become even more difficult.
5 Discussion
Method failure in comparison studies, often manifesting as issues regarding calculation (e.g., non-convergence), runtime, or memory issues, complicates the desired comparison of unconditional performance between the methods under investigation, i.e. across all simulated or benchmark data sets. Our informal literature review and illustrations both make clear that authors should not blindly rely on the initially perceived manifestation of failure but rather make extensive investigations of the sources of failure, which often guides the choice of the handling of the resulting undefined values.
Even if we argue for decisions on a case-by-case basis after careful inspection of the failures, we show and illustrate that some general principles hold in most situations. In particular, method failure usually leads to non-existing values instead of just “missing”, i.e., existent but unobserved values. This means imputation, a common but missing-data-inspired approach to handling method failure, is often inadequate. Another popular and seemingly logical approach, namely discarding data associated with failure (either for all or failing methods), almost never enables the comparison of unconditional performance. In this work, we provide alternative approaches that might be more appropriate to most occurrences of method failure.
An approach that in our view deserves more attention than it receives currently in the literature consists in evaluating whole “pipelines” including one or several fallback strategies in case of method failure. This approach reflects the typical behavior of most users of methods in practice while enabling the comparison of unconditional performance if constructed carefully: if one method does not work, they resort to an alternative one. Future work is required to collect experiences and define standards regarding such pipelines and their evaluation. Above all, transparent reporting of the frequency of failure and of the corresponding handling is always recommended.
A limitation of our recommendations is that, beyond obvious cases, they typically require both high time resources and thorough expertise in the underlying modelling processes and methods. Researchers conducting comparison studies may not always have these resources to the full extent. This particularly applies to a nuanced examination of the interactions between methods and data sets leading to method failure in different instances. We understand that this bottleneck may impair the full implementation of our recommendations in practice, hence our attempt to provide a slimmed-down version of the recommendations. When time resources and expertise are restricted, transparent reporting is all the more important. Also, our viewpoint on method failure as leading to “undefined” performance values does not always hold unrestrictedly, which can complicate the implementation of our recommendations. However, with method failure being a complex matter in and of itself, this underscores the importance of a case-by-case evaluation, as described above.
In this article, method failure has been treated as black-or-white. However, in practice, whether or not a method has failed is itself a value of judgment. One such example is near-separation (van Smeden et al., , 2016), which does not necessarily make itself obvious, hence the existence of dedicated methods to detect it (Mansournia et al., , 2017). Such grey-area failures add further complexities to benchmark and simulation studies: overall performance may depend on the specific criteria used to judge failure.
To sum up, we believe there is not “one” right way to handle method failure in all situations. Instead, it is important to carefully choose an approach depending on the context, and then report, explain (and ideally discuss) the decisions made. With our work, we hope to contribute meaningful ideas on how to think about method failure. They lead to helpful strategies for dealing with it, as outlined as part of our recommendations. However, we also believe that undefined performance values are only one of many issues related to comparison studies that deserve much more time, efforts and methodological strength towards more reliable evidence on the behavior of methods in the context of a “replication crisis in methodological research” (Boulesteix et al., , 2020). We need more well-designed neutral comparison studies (Boulesteix et al., , 2017)—late phase studies according to the terminology of Heinze et al., (2024). In such studies, no time is spent on the development of new methods. This leaves time for implementing our recommendations and diving into the technical details of the methods under investigation.
Funding Information
This work is supported in part by funds from the German Research Foundation (DFG: BO3139/7 and BO3139/9-1). TPM was funded by the UK Medical Research Council (grant MC_UU_00004/09).
Acknowledgements
The authors thank F. Julian D. Lange for useful literature input and Clemens Kreutz for valuable comments. The authors further thank Savanna Ratky for language corrections and Luzia Hanßum for support with formatting the manuscript.
Conflicting interests
The authors have declared no conflicts of interest for this article.
Data availability statement
The code to reproduce all results from Section 4 are available on Github.

### v2

@@
-On the handling of method failure in comparison studies
+Rethinking the handling of method failure in comparison studies
 Abstract
-Comparison studies in methodological research are intended to compare methods in an evidence-based manner, offering guidance to data analysts to select a suitable method for their application.
+Comparison studies in methodological research are intended to compare methods in an evidence-based manner to help data analysts select a suitable method for their application.
 To provide trustworthy evidence, they must be carefully designed, implemented, and reported, especially given the many decisions made in planning and running.
@@
 Despite an increasing emphasis on this topic in recent literature (focusing on non-convergence as a common manifestation), there is little guidance on proper handling and interpretation, and reporting of the chosen approach is often neglected.
-This paper aims to fill this gap and provides practical guidance for handling method failure in comparison studies.
-In particular, we show that the popular approaches of discarding data sets yielding failure (either for all or the failing methods only) and imputing are inappropriate in most cases.
-We also discuss how method failure in published comparison studies—in various contexts from classical statistics and predictive modeling—may manifest differently, but is often caused by a complex interplay of several aspects.
-Building on this, we provide recommendations derived from realistic considerations on suitable fallbacks when encountering method failure, hence avoiding the need for discarding data sets or imputation.
-Finally, we illustrate our recommendations and the dangers of inadequate handling of method failure through two illustrative comparison studies.
+This paper aims to fill this gap and offers practical guidance on handling method failure in comparison studies.
+After exploring common handlings across various published comparison studies from classical statistics and predictive modeling, we show that the popular approaches of discarding data sets yielding failure (either for all or the failing methods only) and imputing are inappropriate in most cases.
+We then recommend a different perspective on method failure—viewing it as the result of a complex interplay of several factors rather than just its manifestation.
+Building on this, we provide recommendations on more adequate handlings of method failure derived from realistic considerations.
+In particular, we propose considering fallback strategies that directly reflect the behavior of real-world users.
+Finally, we illustrate our recommendations and the dangers of inadequate handling of method failure through two exemplary comparison studies.
 1 Introduction
@@
 These issues have gained increasing attention in recent years, both in the context of simulation studies based on artificially generated data (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024) and in the context of real data-based benchmarking studies (Boulesteix et al., , 2017; Friedrich and Friede, , 2024).
-In this work, we focus on a particular aspect of the design of comparison studies that affects both simulation and benchmark studies, namely “method failure”.
-Method failure occurs when a method under investigation fails to produce an output for a (real or simulated) data set considered in the comparison study.
+In this work, we focus on a particular aspect of the design of comparison studies that affects both simulation and benchmark studies, namely “method failure.” Method failure occurs when a method under investigation fails to produce an output for a (real or simulated) data set considered in the comparison study.
 The issue of method failure in comparison studies has begun to attract attention in recent years (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024), often focusing on non-convergence.
 However, there is little guidance on how to handle method failure appropriately, whether it occurs as non-convergence or in another form.
-This is illustrated by our review of published comparison studies from classical statistics and predictive modeling in Section 2 which shows that (reported) handlings often differ greatly between comparison studies.
-Nevertheless, these studies are positive examples of the reporting and the handling of method failure as this issue often goes unreported in many other comparison studies.
-For instance, in a systematic review of $42$ published simulation studies on the analysis of complex longitudinal patient-reported outcomes data, less than half “acknowledge that model non-convergence might occur” (Hinds et al., , 2018).
-Even more notably, only $12$ of $85$ applicable research articles reviewed by Morris et al., (2019) reported convergence as a performance measure.
+In particular, published reviews have shown that method failure goes unreported in the majority of cases.
+For instance, in a systematic review of $42$ published simulation studies on the analysis of complex longitudinal patient-reported outcomes data, less than half “acknowledge that model non-convergence might occur(Hinds et al., , 2018).
+Even more notably, only $12$ of $85$ applicable research articles reviewed by Morris et al., (2019) report convergence as a performance measure.
 This contrasts with our perception—both through our own experiences and informal discussions with colleagues—that the problem of method failure is highly prevalent.
-In particular, signs of method failure are often traceable in the code provided for reproducibility even when not reported in the manuscript.
-This is often clear through the use of R functions try() or tryCatch() for the handling of errors, for example.
-To illustrate the difficulty of handling method failure adequately, imagine conducting a comparison study in statistics or predictive modeling to compare the performance of a selection of methods.
-In the context of classical statistics, you likely focus on comparing the methods’ performances in “estimating one or more population quantities” (Morris et al., , 2019), sometimes called “estimand(s)” and base your analysis on data simulated using a certain data-generating mechanism (DGM).
-In predictive modeling, you are likely to compare the methods’ predictive performances, that is, their ability to make predictions on a new set of observations.
+In particular, signs of method failure are often traceable in the code provided for reproducibility, even when not reported in the manuscript.
+This is often clear through the use of explicit error-handling mechanisms (e.g., functions try() or tryCatch() in R or try-except constructs in Python).
+To illustrate the difficulty of handling method failure adequately, imagine a researcher conducting a comparison study in statistics or predictive modeling to compare the performance of three methods based on four simulation/benchmark data sets.
+In the context of classical statistics, they likely focus on comparing the methods’ performances (e.g., bias) in “estimating one or more population quantities” (Morris et al., , 2019), sometimes called “estimand(s),” and base the analysis on data simulated using a certain data-generating mechanism (DGM).
+In predictive modeling, they are likely to compare the methods’ predictive performances (e.g., accuracy), that is, their ability to make predictions on a new set of observations.
 Here, the analysis is typically performed on real benchmarking data sets.
-For simplicity, we refer to both settings as “simulation studies” and “benchmark studies”, respectively (though admitting that there may be intermediate situations, e.g.
-simulation studies investigating predictive performance on artificial data).
-During your experiments, you encounter the failure of a method in one or more of the following ways: (i) you receive an error message or output of NA (“not applicable”) or NaN (“not a number”), implying that a certain calculation could not be completed (e.g., non-convergence), (ii) your system crashes, or (iii) the experiment runs for an excessively long time, forcing you to abort.
-Table 1: Overview table of (a) a fictive simulation study and (b) a fictive benchmark study encountering method failure.
-This results in undefined estimates values for the simulation study (a) and undefined performance values for the benchmark study (b), marked in blue.
-In both cases, this complicates the aggregation of performance of the given method.
-(a) Fictive simulation study: overview of estimates produced by three methods under investigation (columns) across $4$ repetitions, i.e.
-simulated data sets (rows), in one simulation scenario, where the true value is equal to $4$.
-An estimates value is obtained for each of the three methods and four simulation repetitions.
-Bias is commonly calculated as the mean deviation of the estimates from the true value across the $4$ repetitions.
-Method failure first leads to an undefined estimates value in the given repetition, which then complicates the estimation of performance of the affected method.
+For simplicity, we refer to both settings as “simulation studies” and “benchmark studies,” respectively (though admitting that there may be intermediate situations, e.g., simulation studies investigating predictive performance on artificial data).
+During the experiments, the researcher encounters the failure of some methods in one or more of the following ways: (i) they receive an error message or output of NA (“not applicable”) or NaN (“not a number”), implying that a certain calculation could not be completed (e.g., non-convergence), (ii) their system crashes, or (iii) the experiment runs for an excessively long time, forcing the researcher to abort.
+While the majority of real published comparison studies do not report method failure, there are positive examples from classical statistics and predictive modeling in which the occurrence (and oftentimes also the applied handling) of method failure is described.
+For instance, van Smeden et al., (2016) investigate why different simulation studies offer conflicting minimal “events per variable (EPV)” recommendations for logistic regression.
+All three studies featured, exploring the effect of EPV on performance measures such as bias and coverage, encounter separated data sets (i.e., outcome variable is perfectly predicted by one or a combination of multiple covariates), which lead to convergence issues of the maximum likelihood process.
+That is, convergence is not achieved within the pre-specified number of iterations, or the process converges to a point that is not the maximum likelihood estimate (van Smeden et al., , 2016).
+In a study from the statistical learning field, Gijsbers et al., (2024) conduct a comparison study of nine automated machine learning methods (“AutoML frameworks”).
+They encounter and report method failure in the forms of (i) exceeding available memory (or other memory-related issues), (ii) exceeding time limits, (iii) failure related to data set characteristics (such as highly imbalanced data), and (iv) failure caused by errors in the implementation of the AutoML method (Gijsbers et al., , 2024).
+Table 1: Overview table of (a) a fictive simulation study across four repetitions (rows) and (b) a fictive benchmark study across four benchmark data sets (rows) for three methods (columns).
+Method failure results in undefined values for the affected combination of data set and method, marked by “NA.”
+(a) Fictive simulation study: overview of estimates, with the true value being equal to $4$.
+Method failure first leads to an undefined estimates value for the given method–data set combination, complicating the derivation of bias for that method.
 Repetition Method 1 Method 2 Method 3
@@
 Bias ? 0.07 ?
-(b) Fictive benchmark study in predictive modeling: an estimated accuracy is obtained for each of the four data sets (rows) and three methods under investigation (columns).
-Method failure directly leads to an undefined performance (i.e., accuracy) value for the given method in the given row of the table, complicating the aggregation of its performance over all rows required to obtain an overall performance estimation.
-Benchmark data set Method 1 Method 2 Method 3
+(b) Fictive benchmark study: overview of estimated accuracies across all method–data set combinations.
+Method failure directly leads to an undefined performance (i.e., accuracy) value for the given method–data set combination.
+Benchm. data set Method 1 Method 2 Method 3
 1 0.85 0.88 0.87
@@
 Average accuracy ? 0.84 ?
-Table 1 depicts a hypothetical summary of the results of your study in each simulation repetition/benchmark data set, denoting the undefined values by a blue “NA”.
-You may now wonder how to accurately determine the methods’ performances from this incomplete information.
-Note that Table 1 first of all highlights an important distinction that must be made between simulation and benchmark studies regarding the consequences of method failure.
-In a simulation study focusing on assessing the methods’ performances in estimating an estimand, a method typically produces a single estimate for the entire data set in the given repetition (i.e., row in Table 1(a)).
-Its performance is estimated as an aggregate over all rows in Table 1(a).
-In this setting, method failure first leads to an undefined estimate value which then complicates the assessment of the affected method’s performance.
-In a benchmark study focusing on the predictive performance of methods, on the other hand, a method makes a prediction for each row of the corresponding test data for a given benchmark data set (based on the decision rule generated on the training data).
-The method’s performance (e.g., accuracy) is then measured directly in each real data set (i.e., each row in Table 1(b)) and its overall performance is averaged over all rows of Table 1(b).
-Method failure results directly in an undefined performance value.
-Looking at Table 1, it is tempting to view method failure as missing data: in each case, the row summarising results contains “holes” where results were expected.
-We caution against this.
-Missing values are data that exist and were intended to be recorded but for one reason or another are “not observed” (Rubin, , 2004; Carpenter et al., , 2023).
-In our context, although we intended to record a value, method failure means that the value does not exist for that particular data set.
-This is similar to data truncation by death as opposed to the inability to collect data due to the non-appearance of a patient at the doctor’s office (missing data).
-As a result, some of the most commonly used approaches to handling method failure such as imputation, being “missing data inspired”, are inadequate in most cases.
-Furthermore, only few are aware that the typical ways and reportings of discarding data sets associated with failure (either for all or the failing methods only) do not enable meaningful method comparisons.
-We illustrate this in detail in Sections 3 and 4.
-In the subsequent part of this paper, we use the term “undefined” instead of “missing” to emphasize the important distinction between method failure and the “regular” missing data context.
-The consequence of our dissuasion from “missing data-inspired” approaches is that our work focuses intensively on alternative ways to deal with method failure.
+Regardless of the form in which a researcher encounters method failure, it leads to NA for the associated method-data set combination where otherwise an estimate (for simulation studies) or a performance value (for benchmark studies) is expected.
+Table 1 depicts a corresponding summary of the results of the hypothetical researcher’s study across all simulation repetition/benchmark data sets and methods.
+In simulation studies, the NAs complicate the derivation of performance (e.g., bias as the mean deviation of the estimates from the true value across all repetitions).
+In benchmark studies, on the other hand, they hinder the aggregation of performance values into an average performance (e.g., average estimated accuracy as the mean fraction of correct predictions across all benchmark data sets).
+In both cases, method failure complicates the assessment and especially the comparison of performance of the methods under investigation such that the researcher finds themself wondering how to proceed adequately.
+The above-shown examples of published comparison studies alone already illustrate different viewpoints on method failure and correspondingly different ways in which it is handled in practice.
+van Smeden et al., (2016) observe that separated data sets are commonly removed from the analysis.
+This, however, affects bias, mean squared error (MSE), and confidence interval width, even if the frequency of separated data sets is small (van Smeden et al., , 2016).
+Gijsbers et al., (2024) argue against discarding data by describing that method failure is typically correlated with data set characteristics.
+Instead, they impute by the performance of a constant predictor.
+Examining additional published studies, which we do throughout this paper, shows that imputation and discarding data sets associated with failure are among the most common practices.
+However, it also becomes clear that they are only two of many handlings applied in practice and that others, such as modifying the affected methods’ implementations and changing software, are common as well—an indicator that proper guidance is commonly lacking.
+Our paper aims to fill this gap and provides detailed instructions on how to handle method failure.
+In particular, we illustrate why usual handlings of method failure derived from treating the resulting NAs as regular “missing data,” and especially imputation and discarding data, are inappropriate in most settings (leading us to refer to “undefined” rather than “missing” performance values in instances of method failure).
+As an alternative, we promote considering method failure from a different viewpoint, namely as the result of a complex and individual interplay of multiple factors—which we use to derive alternative and more appropriate measures that can be applied when encountering method failure.
+Our paper targets both the planning and the execution stage of a comparison study.
+While many instances of method failure can only be directly observed—and therefore investigated—after the study is executed, certain considerations should be made while planning to anticipate and ideally reduce the potential for failure.
 The rest of this paper is structured as follows.
-In Section 2, we review and summarize the different manifestations of method failure in published comparison studies.
-In particular, we discuss how at the core, they often result from a complex interplay of multiple factors, and that an adequate handling requires thorough investigation on a case-by-case basis.
-Building on this, Section 3 is a compilation of practical recommendations for handling method failure which differentiates between different scopes of the study and research stages in which it may take place.
-In particular, we promote using fallback strategies when encountering method failure.
-This approach, if based on thorough and realistic considerations, enables the aggregation of performance across data sets even when method failure occurs.
-However, considering possible time constraints, we also discuss “slimmed-down” ways in which authors can still implement our recommendations.
+In Section 2, we give an overview of the status quo by reviewing and summarizing the different manifestations and handlings of method failure reported in published comparison studies, followed by a detailed discussion of why discarding data or imputing values are usually inadequate when encountering method failure.
+Building on this, Section 3 is a compilation of alternative recommendations for handling method failure, which differentiates between different scopes of the study and research stages in which it may take place.
+In particular, we promote using fallback strategies when encountering method failure, which enables the aggregation of performance across data sets even when method failure occurs and on top of that, reflects the behavior of real method users across many settings.
+Considering possible time constraints, we also discuss “slimmed-down” ways in which authors can implement part of our recommendations if unable to follow them completely.
 Finally, in Section 4, we demonstrate our recommendations using two fictive comparison studies that illustrate the complexity of the discussed issues.
-2 Method failure in the literature
-In our own experience, method failure occurs frequently when conducting comparison studies in statistics and predictive modeling.
-As already described, a look at the literature suggests that others also encounter this problem, either because method failure is explicitly mentioned in the text or because signs of method failure can be found in the code provided by the authors for reproducibility.
-Section 2.1 briefly outlines how method failure manifests and is handled in a selection of published studies from the classical statistics and predictive modeling literature.
-In Section 2.2, we argue that at the core, method failure is a complex interplay of multiple factors, deserving particular attention.
-Finally, Section 2.3 briefly categorizes different procedures for handling method failure used in these studies and provides a first outlook on the recommendations provided in Section 3.
-2.1 Reporting and handling of method failure from the literature: vignettes
-The following comparison studies should be viewed as positive examples, being among the few that report the occurrence of method failure, often alongside reasons and strategies for handling the resulting undefined values when summarizing and interpreting the results.
-The strategies adopted in these studies inspire—but do not completely overlap with—our recommendations given in Section 3.
-In these cases, we do not intend to criticize the applied approaches but aim to further improve the handling and reporting of method failure for future work.
-van Smeden et al., (2016) investigate why different simulation studies offer conflicting minimal “events per variable” recommendations for logistic regression.
-They particularly observe that separated data sets (i.e., outcome variable is perfectly predicted by one or a combination of multiple covariates) are commonly omitted from the analysis since they lead to convergence issues of the likelihood maximization process.
-However, van Smeden et al., (2016) demonstrate that the removal of these data sets, in which the association between the covariates and outcomes is particularly strong, substantially affects bias, mean squared error (MSE), and confidence interval width, even if the frequency of separated data sets is small.
-In Voeltz, (2024), a simulation study is conducted to compare three frequentist methods for the meta-analysis of diagnostic accuracy studies regarding sensitivity and specificity.
-Two of the three methods under investigation are occasionally affected by method failure due to non-convergence (while the third one does not employ an iterative process).
-In the case of non-convergence, one of the two does not return an error but outputs the current, non-converged estimate.
-As these outputs are often used in practice nevertheless, the authors also include them in the performance comparison.
-For the remaining method affected by non-convergence, they discard the data sets for which it fails and use the proportion of simulation repetitions in which the methods converge as an additional performance measure.
-An extreme example of excluding performance data is to remove an entire method.
-This is done by Ruxton and Neuhäuser, (2013) in a simulation study to compare the empirical coverage of methods commonly used to calculate confidence intervals for the odds ratio.
-Given a binary outcome and binary exposure variable, some methods fail due to division by 00 if the corresponding $2\times 2$ contingency table contains one or more zero-entries.
-Masaoud and Stryhn, (2010) compare the performance of different marginal and random effects estimation methods for the analysis of binary longitudinal data (e.g., GEE, Bayesian Markov Chain Monte Carlo) in a simulation study.
-With data simulated using two different data-generating mechanisms (i.e., stemming from marginal and random effects models), methods are compared regarding bias and coverage of the confidence intervals (as well as efficiency for data simulated from the marginal model).
-The authors encounter instances of “non-convergence or non-sensible estimates” (Masaoud and Stryhn, , 2010), for which the handling varies for the different methods affected.
-They either (i) change the software from R to SAS, (ii) switch the optimization technique used internally by the method, or (iii) for GEE, resort to using model-based standard errors if the method fails to produce meaningful robust standard errors.
-However, if method failure cannot be solved using these measures, they discard the corresponding data set.
-Dunias et al., (2024) conduct a simulation study to compare the performance of four hyperparameter tuning procedures (including $k$-fold cross validation and the one-standard error rule for tuning) for prediction methods developed using predictive modeling methods such as Ridge, LASSO, or Random Forest.
-For some data sets, LASSO does not select any predictor variable, leading to a constant predictor.
-Resulting in the inability to assess the performance in terms of the calibration slope, the “[missing] calibration slopes [are] replaced by the maximum value of calibration slope for the corresponding scenario” (Dunias et al., , 2024).
-Note that this corresponds to imputation by the worst “observable” value since higher values of the calibration slope generally indicate that “risk predictions are too narrowly distributed” (Dunias et al., , 2024).
-An approach to handling method failure by imputation is applied by Bischl et al., (2013) in their comparison study of $11$ methods for classification on $49$ real and simulated (benchmarking) data sets.
-They use misclassification error as performance measure combined with $100$ subsampling (also called “Monte Carlo cross-validation”) repetitions.
-When encountering method failure, they impute based on the proportion of repetitions in which the affected method fails.
-If below $20\%$, they impute by sampling from a normal distribution that is estimated from the remaining performance values of that model.
-This way of imputing serves to adhere to the assumptions of the model used to test the generalization performance.
-If a method fails for more than $20\%$ of repetitions, they impute by the worst possible value with the reasoning that the method’s behavior is “too unreliable for the current data set” (Bischl et al., , 2013).
-Herrmann et al., (2021) perform a benchmark study to compare the predictive performance of survival time prediction methods based on high-dimensional multi-omics data using real cancer data sets.
-They encounter method failure in several cross-validation folds for several methods (including one method with excessive runtimes).
-Inspired by Bischl et al., (2013), they impute undefined performance values based on the proportion of “failed” cross-validation folds.
-If it is above $20\%$, they impute by the “worst possible value” of the corresponding metric, else, they impute by the mean performance of the method in the remaining “successful” folds.
-Note that the latter is equivalent to discarding those data sets in the performance assessment for which the given method fails.
-In a subsequent project, Nießl et al., (2022) investigate the effect of different imputation procedures on the performance assessment and ranking of the methods.
-They consider alternative procedures including imputation by the worst possible value and by a value that weights the method’s overall proportion of undefined performance values and its performance in the remaining folds.
+2 Method failure in published comparison studies
+In Section 2.1, we summarize common manifestations and handlings of method failure reported in the literature.
+Since some authors focus on how method failure is handled rather than explaining how it occurred, the corresponding studies are discussed only in Section 2.1.2.
+An overview of all study contexts, manifestations, and handlings of method failure are provided in Table 2.
+2.1 Status quo: occurrence and handling of method failure
+Comparison study     Manifestations     Handlings
+        Resolve Summarize despite
+        method failure method failure
+Author Context of the study     non-convergence (and others) runtime issues
+memory issues
+Remove method Modify implementation Change software Discard data for failing methods Discard data for all methods Impute Performance measure
+van Smeden et al., (2016)a Simulation study to investigate why different published simulation studies offer conflicting minimal “events per variable (EPV)” recommendations for logistic regression.
+Features three published simulation studies exploring the effect of EPV on performance measures.
+$\times$     $\times$
+Gijsbers et al., (2024) Comparison study of nine automated machine learning (“AutoML”) methods.     $\times$ $\times$
+$\times$
+$\times$
+Zapf et al., (2024) Simulation study to compare three frequentist methods for the meta-analysis of diagnostic accuracy studies regarding sensitivity and specificity.     $\times$     $\times$ $\times$
+Ruxton and Neuhäuser, (2013) Simulation study on the empirical coverage of eight methods for calculating confidence intervals for the odds ratio for binary outcome and exposure.
+$\times$     $\times$
+Hornung et al., (2024) Benchmark study using thirteen multi-omics data sets to assess the performance of seven prediction methods for multi-omics data that can handle block-wise missingness by artificially inducing specific patterns of block-wise missing values into initially complete data.
+$\times$
+$\times$
+$\times$
+Masaoud and Stryhn, (2010) Two simulation studies to compare the performance of seven marginal and random effects estimation methods for binary longitudinal data (e.g., GEE, Bayesian Markov Chain Monte Carlo).
+$\times$     $\times$ $\times$ $\times$
+Crowther et al., (2014) Simulation study to assess the performance of a multilevel mixed effects parametric survival model using adaptive and nonadaptive Gauss-Hermite quadrature.
+$\times$     $\times$
+Dunias et al., (2024) Simulation study to compare the performance of five hyperparameter tuning procedures for clinical prediction models built upon statistical learning methods.
+$\times$     $\times$
+Fernández-Delgado et al., (2014) Benchmark study of 179 supervised classification methods across 121 benchmark data sets.     $\times$
+$\times$
+$\times$ $\times$
+Herrmann et al., (2021) Benchmark study on the predictive performance of thirteen survival time prediction methods based on eighteen high-dimensional multi-omics data.
+* $\times$ *     $\times$ $\times$
+Bischl et al., (2013) Comparison study of eleven classification methods on 49 real benchmark and simulated data sets.     * * *     $\times$ $\times$
+Nießl et al., (2022)b Meta-study to investigate the effect of different imputation procedures on the performance assessment and ranking of the methods.
+Considered alternative procedures include imputation by the worst possible value and by a value that weighs the method’s overall proportion of method failure against its performance in the remaining folds.
+* * *     $\times$
+  • a
+    van Smeden et al van Smeden et al., (2016) observe that separated data sets (i.e., data sets leading to non-convergence) are commonly omitted from analyses.
+  • b
+    Nießl et alNießl et al., (2022) perform an example benchmark study based on the results of the study by Herrmann et alHerrmann et al., (2021) to illustrate how different ways of imputation can influence the results of a benchmark study.
+    The authors did not report details on the manifestation of method failure.
+Table 2: Selection of comparison studies from statistics and predictive modeling that report the occurrence and handling of method failure.
+2.1.1 Common manifestations of method failure
+Non-convergence and other calculation issues  Authors often encounter the inability of a method to perform a calculation necessary for prediction or estimation, either implied through an error, a warning, or a meaningless output.
+For methods employing an iterative procedure (e.g., an iterative maximum-likelihood procedure), this is called “non-convergence,” i.e., the method fails to produce a valid output within the pre-specified number of iterations.
+Examples  Zapf et al., (2024) encounter non-convergence for two of the three methods under investigation.
+One returns an error, while the other outputs the current, non-converged estimate in these instances (Zapf et al., , 2024).
+Masaoud and Stryhn, (2010) encounter instances of “non-convergence or non-sensible estimates” (Masaoud and Stryhn, , 2010) for some method-data set combinations.
+In the study by Crowther et al., (2014), the methods are based on numerical integration and encounter estimation difficulties (non-convergence) with “challenging data sets(Crowther et al., , 2014).
+Hornung et al., (2024) report instances of method failure when methods are “not applicable” (Hornung et al., , 2024) to certain differences between training and test sets regarding the patterns of block-wise missingness.
+Additionally, they report calculation issues for one method explicitly disallowing certain structures in the training data set, resulting in failure.
+Calculation issues are reported by Ruxton and Neuhäuser, (2013), where some odds ratio estimation methods fail due to division by 00 when the $2\times 2$ contingency table contains one or more zero-entries.
+Another case of calculation issues can be found in the simulation study conducted by Dunias et al., (2024).
+For some data sets, the statistical learning method LASSO does not select any predictor variable, leading to a constant predictor and preventing the assessment of performance regarding calibration slope.
+Memory issues  Even before exceeding possible time budgets, some methods crash due to memory issues. That is, the given method consumes more memory than is available.
+Examples  Gijsbers et al., (2024) report memory constraint violations, segmentation faults, and Java server crashes as instances of method failure related to memory.
+Hornung et al., (2024) also encounter memory issues, which they attribute to a specific missingness pattern induced into the training data.
+Fernández-Delgado et al., (2014) encounter memory issues in combination with large data sets.
+Runtime issues  Another form of method failure often reported in comparison studies is excessive runtime.
+Some studies set fixed time budgets, so exceeding them is considered method failure.
+Others without fixed budgets encounter method failure when experiments for single methods run for an excessive amount of time (e.g.
+days), forcing the researcher to abort.
+Unlike non-convergence and memory issues, method failure is actively declared by the researchers when runtime issues occur, as without these (naturally necessary) limitations, a method might eventually generate an output at some point.
+Resulting implications are explored in Section 3.1.6.
+Examples  While Gijsbers et al., (2024) encounter the excess of predefined time budgets for model training, Herrmann et al., (2021) face runtime issues without fixed budgets.
+They encounter “computations lasting several days for one single model fit for large data sets” (Herrmann et al., , 2021) for one method.
+2.1.2 Common approaches to handling method failure
+Based on the published comparison studies, the handling of method failure can be commonly divided into two general approaches.
+Note that a literature review by Pawel et al., (2024), yielding a structured investigation of the frequencies of handling approaches from published simulation studies, results in four main handlings, which are among those we have commonly observed.
+Resolving method failure  Some authors resolve method failure completely so that no instances remain.
+A straightforward way observable in the literature is to completely remove the affected method from the analysis.
+For instance, Ruxton and Neuhäuser, (2013) exclude all methods that fail with zeros in the $2\times 2$ contingency table.
+Masaoud and Stryhn, (2010) use method removal as a last resort in their sequence of handlings when previous measures (see below) have been unsuccessful.
+This illustrates well that oftentimes, different handlings are applied within the same study, sometimes switching between approaches or applying them sequentially until successful.
+Note that a more moderate form of method removal can be observed in the benchmark study by Herrmann et al., (2021).
+For the method encountering runtime issues, they restrict the analysis to the subset of smaller benchmark data sets and report the corresponding results separately from the remaining methods.
+An alternative to removing methods is to modify their implementation such that an output can be successfully produced.
+For instance, in the study by Masaoud and Stryhn, (2010), “different optimization techniques were tried” (Masaoud and Stryhn, , 2010) for one method when the default procedure resulted in failure.
+Crowther et al., (2014) react to method failure by successively increasing the number of quadrature points until convergence is reached.
+Lastly, changing software for the method affected by method failure can also be observed in the literature.
+Switching the operating system can be observed in the study by Herrmann et al., (2021), who for one method encounter “a fatal error in R under Windows, but not using the Linux distribution Ubuntu 14.04” (Herrmann et al., , 2021).
+Masaoud and Stryhn, (2010), while remaining in the same operating system, generate the required output for one method in SAS when it fails in R.
+Summarizing performance despite method failure Instead of resolving method failure, some authors summarize the methods’ performances despite existing failures.
+A widespread approach is to discard data sets associated with failure.
+The first version of this approach, consisting of discarding data sets for the failing methods only, is followed by Zapf et al., (2024) for the method affected by errors due to non-convergence.
+Fernández-Delgado et al., (2014), comparing the methods’ performance based on two performance measures, discard the associated data sets for the methods affected by method failure in the comparison based on average accuracy.
+The second version of discarding data, namely discarding the corresponding data sets for all methods, is applied by Hornung et al., (2024).
+Particularly, they argue against discarding the data sets for the failing methods only by saying that it would not yield a fair comparison since predictive performance is generally associated with the induced pattern of block-wise missingness (Hornung et al., , 2024).
+Instead of discarding data, it is also common to impute performance values for the instances of method failure.
+Note that imputation yields a variety of options.
+Fernández-Delgado et al., (2014), for instance, impute by the mean accuracy of all remaining methods on the given data set for their second performance measure (Friedman ranking) when a method fails as it requires the same number of performance values across all methods.
+On the other hand, recall Gijsbers et al., (2024) and their comparison study of AutoML methods in which they impute with a constant predictor when a method fails.
+They choose this “very penalizing imputation strategy(Gijsbers et al., , 2024), which can be viewed as a form of imputation by the worst possible value, with a similar reasoning to Hornung et al., (2024).
+That is, failure may be correlated with data set characteristics such that discarding data sets for the failing methods or imputing by the average performance of the given method on the remaining data sets may unfairly favor the failing methods (Gijsbers et al., , 2024).
+Note that Dunias et al., (2024) also impute by a constant predictor when a statistical learning method fails.
+An alternative way of imputing is employed by Bischl et al., (2013), who impute performance values based on the proportion of repetitions in which the affected method fails.
+Below $20\%$, they impute by sampling from a normal distribution estimated from the remaining performance values of that model.
+Above $20\%$, on the other hand, they impute by the worst possible value with the reasoning that the method’s behavior is “too unreliable for the current data set” (Bischl et al., , 2013).
+A similar imputation strategy is employed by Herrmann et al., (2021).
+While using the same imputation rule when a method fails on more than $20\%$ of data sets, they otherwise impute by the mean performance of the method in the remaining “successful” data sets.
+In a subsequent project, Nießl et al., (2022) consider alternative procedures, including imputation by the worst possible value and by a value that weights the method’s overall proportion of method failure against its performance in the remaining folds.
 They show that the approach to imputing the undefined performance values has a notable effect on the performance ranking between the methods.
-Another example can be found in the work of Hornung et al., (2024).
-They assess the performance of prediction methods that can handle block-wise multi-omics data by artificially inducing specific patterns of block-wise missing values into initially complete multi-omics data.
-Three of the considered methods fail for some data sets (with a maximum proportion of $34\%$).
-In some rare cases, failure is caused by memory issues.
-More commonly, however, the methods are “not applicable” (Hornung et al., , 2024) if the training and test data sets differ in the induced pattern of missing data blocks in certain ways.
-The resulting undefined performance values are addressed by discarding the corresponding data sets for all methods with the reasoning that discarding them only for the failing methods would lead to confounding by the induced pattern of block-wise missing values.
-Note that Hornung et al., (2024) argue that instances of memory issues are likely caused by a specific missingness pattern induced into the training data.
-Fernández-Delgado et al., (2014) perform an extensive comparison study of supervised classification methods (“classifiers”) from several model families.
-The classifiers are assessed in $121$ (real) benchmarking data sets in combination with $4$-fold cross-validation and “accuracy” as performance measure (which is then averaged over all benchmarking data sets).
-Additionally, they compute the Friedman ranking (Sheskin, , 2011) based on accuracy.
-They encounter method failure in $2.1\%$ of classifier-data set combinations.
-As reasons for failure, they list “collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al., , 2014).
-Note that equal predictor values of all observations in the training data lead to the failure of the corresponding method since this violates its requirements.
-Memory issues also lead to method failure with large data sets, while particularly low numbers of observations in some classes in small data sets lead to errors for some classifiers.
-The authors discard the corresponding data sets for the failing methods to obtain each classifier’s accuracy.
-For the Friedman ranking, which requires an identical number of accuracy values for all classifiers, they impute undefined accuracy values with the mean value across all classifiers for this given data set.
-Gijsbers et al., (2024) introduce a benchmarking tool for automated machine learning (“AutoML”) frameworks, which are considered as the “methods” under comparison.
-They conduct a comparison study of nine such methods based on various data sets using the Area Under the Curve (AUC) as performance metric.
-They categorize the reasons for failure into (i) exceeding available memory (or other memory-related issues), (ii) exceeding time limits, (iii) failure related to data set characteristics (such as highly imbalanced data), and (iv) failure caused by errors in the implementation of the AutoML framework.
-Categories (i) and (ii) are reported as the most common.
-Gijsbers et al., (2024) argue against ignoring undefined performance values, stating that they do not occur at random but are correlated with data set characteristics.
-They impute by the performance of a constant predictor.
-2.2 Typical manifestations of method failure
-Summarizing from the previous Section 2.1, there are common ways in which method failure often reveals itself to authors of comparison studies; see Table 2 for an additional overview.
-Table 2: Common manifestations of method failure
-Manifestation Explanation Examples from literature
-Non-convergence
-and other
-calculation issues
-Method cannot perform required calculations (within maximum number of iterations), therefore fails to produce a stable output. Cannot be solved by increasing memory or time budgets.
-$\rightarrow$ Example: - division by zero - attempt to invert singular matrix $\rightarrow$ Can result in: - error message - non-meaningful output such as NA, NaN, output of $\infty$, or inflated estimation (for estimation task)
-van Smeden et al., (2016)
-Fernández-Delgado et al., (2014)
-Memory
-issues
-Method consumes more memory than available to perform required computations.
-$\rightarrow$ Can result in: system crashes (“fatal error”)
-Hornung et al., (2024)
-Fernández-Delgado et al., (2014)
-Runtime
-issues
-Method fails to perform required computations within pre-specified or feasible time limits.
-$\rightarrow$ “Failure by design” of study
-$\rightarrow$ Real users would likely obtain valid output
-Herrmann et al., (2021)
-Gijsbers et al., (2024)
-Non-convergence and other calculation issues  Authors often encounter the inability of a method to perform a calculation necessary for prediction or estimation, either implied through an error, a warning message, or a method output that is not meaningful.
-For methods employing an iterative procedure (e.g., an iterative maximum-likelihood procedure), this is often referred to as “non-convergence”, i.e., the method fails to produce a stable and valid output in the pre-specified number of iterations.
-Memory issues  Even before exceeding possible time budgets, some methods crash due to memory issues.
-That is, the given method consumes more memory than is available.
-Gijsbers et al., (2024) additionally list segmentation faults as memory-related errors.
-Runtime issues  A variant of method failure manifested and reported in existing comparison studies is excessive runtime.
-Note that in some comparison studies, fixed time budgets are predefined (possibly with an additional leniency period) such that exceeding them is declared method failure (Gijsbers et al., , 2024).
-Other studies do not predefine fixed budgets but find that, due to excessive runtimes, running the entire experiment for the affected method is impossible and must therefore be aborted (Herrmann et al., , 2021).
-Note that in contrast to the other two manifestations, method failure is actively declared by the authors when runtime issues occur.
-That is, running the methods under investigation on hundreds or thousands of data sets in the study creates the necessity of time restrictions, which are set by the authors arbitrarily, and without which the method would likely produce a valid output.
-In particular, in many cases, a corresponding real-world user, running the method on a single data set only, would likely obtain a valid output.
-This has important implications for the handling of method failure, which we will discuss in Section 3.2.
-At first glance, it may seem like calculation, memory, and runtime issues are separate reasons for method failure, such that authors report specific “reasons” for failure for different methods within their study or even different “reasons” for a single method in different instances.
-For instance, some authors report non-convergence as a “reason” for failure in one method–data set combination and memory issues in another.
-However, upon closer examination, it becomes evident that these are not so much reasons but rather manifestations of method failure that are just the tip of the iceberg.
-In essence, method failure (or what might appear as such) is often the result of an individual and complex interplay of many different factors.
-For example, Kriegel et al., (2017) show that runtimes can depend on numerous factors: “You might be surprised by the amount of differences in runtime that we point out, spanning up to several orders of magnitude in different implementations of one and the same algorithm.
-We found major performance differences even within the same toolkit, in different versions or releases: due to bugs, algorithmic improvement (having an indirect impact on the tested algorithm, e.g., in data handling), regressions, and environment changes.
-Finally, it is also noteworthy that good or bad choice of parameters for different algorithms can lead to void runtime comparisons”.
-It is important to emphasize that differences in hardware (e.g., available memory) and/or software (e.g., operating system) can also affect assessment.
-As an extreme example of software dependency, consider that Herrmann et al., (2021) reported that one of the methods under investigation (sparse group lasso) was completely unusable due to an error in R that occurred using Windows as an operating system, while it did not occur under Ubuntu.
-As an extreme example of hardware dependency, consider that the feasibility of deep learning to train large artificial neural networks tremendously increases with the use of graphics processing units (GPUs).
-In essence, this means it is important to keep in mind that a computation that fails or is infeasible in terms of runtime and budget on one specific system setup, may work on a different setup.
-Note that the reverse situation that a computation that can be successfully conducted on one specific system cannot be conducted—i.e., reproduced—using a (slightly) different setup (e.g., during review), is very common and addressed by the term computational reproducibility.
-In addition to these (combinations of) factors associated with a method (or its implementation) and the technical issues mentioned above, failure is typically also related to certain characteristics of the given data.
-That is, some characteristics can complicate required calculations for a method such that they are completely impossible or cannot be completed within the maximum number of iterations (non-convergence), or they exceed available time or memory budgets.
-For instance, van Smeden et al., (2016) report non-convergence of logistic regression algorithms when the outcome variable is perfectly separated by one or more predictor variables.
-To sum up, oftentimes, certain characteristics of the data cause a certain method, in its current implementation, to fail in a certain computational environment.
-2.3 Common approaches to handling method failure
-The approaches to handling method failure from the published comparison studies in Section 2.1 can generally be divided into two categories.
-First, some authors resolve method failure so that no undefined values remain.
-This includes entirely removing failing methods from the study or modifying its implementation (e.g., adjusting some parameters) to make it work for all data sets.
-Second, other authors summarize the affected method(s) performance(s) despite undefined performance values.
-Most commonly, the corresponding data sets are discarded either for all or only for the failing methods.
-Other authors impute undefined performance values.
-Again, this is often done in various ways, for example, imputation through the worst possible value (when such a value exists) or through the worst (or mean) value achieved by other methods on the particular data set.
-Additionally, some authors also report the proportion of (real or simulated) data sets for which each method fails.
-In the following Section 3, we discuss why imputation and discarding data (either for all or the failing methods only) are inadequate in most cases of method failure.
-In particular, we recommend alternative and more appropriate approaches that are based on realistic considerations regarding the research stage in which the study takes place and the corresponding targeted users.
-3 Principles and recommendations for handling method failure
-3.1 Absolute, relative, unconditional and conditional performance
-There is no one-size-fits-all solution to handling method failure, exemplified by the examples in Section 2.
+When summarizing performance despite the failure of an existing method, some authors report the proportion of (real or simulated) data sets for which each method fails.
+For instance, Hornung et al., (2024) list “the frequency of data sets with missing results” (Hornung et al., , 2024) for the methods affected by method failure.
+Gijsbers et al., (2024) provide multiple failure plots, consisting of the number of failures per method with differentiation of the different failure types, the number of failures by size of the associated data set, and boxplots of the training durations for each method with indicators of excessive runtimes.
+Zapf et al., (2024) even include convergence as a performance measure, supported by a table of the number of converged simulation runs for each simulation scenario across the methods in the supplement.
+2.2 Why popular handlings are often inadequate
+2.2.1 Why discarding data is inadequate
+Currently, there is no one-size-fits-all solution to handling method failure, exemplified by the examples in Section 2.
 This section discusses some principles to guide decisions.
@@
 relative performance.
-Unconditional performance simply means “in all data sets”, while conditional means “in data sets with no method failure”.
-The unconditional bias of method B is 0.14, while the unconditional bias of method A is not defined.
+Unconditional performance simply means “in all data sets,” while conditional means “in data sets where this method did not fail.” The unconditional bias of method B is 0.14, while the unconditional bias of method A is not defined.
 The conditional bias of both methods is 0.09.
 Parallel descriptions apply to the hypothetical benchmark study in Table 3(b) (b).
-Absolute performance means “considered on its own” and relative performance means “when compared with another method”.
-In both Tables 3(b) (a) and 3(b) (b) the relative unconditional performance of methods cannot be quantified because the unconditional performance of method A is undefined.
-The relative conditional performance can be quantified: in Table 3(b) (a), the two methods have the same bias conditional on non-failure of method A; in Table 3(b) (b), the two methods have the same accuracy conditional on non-failure of method A.
+Absolute performance means “considered on its own” and relative performance means “when compared with another method.”
 These terms help to distill the problems with method failure in comparison studies.
-Firstly, the absolute conditional performance of method B should be of little interest.
-In deciding to use method B, why should we care about its performance when method A, a method we are not going to use, does not fail?
-We should care about a method’s unconditional performance.
-The problem is that we cannot compare the methods in terms of their unconditional performance.
-Comparisons should be performed vertically, or we are not comparing like with like: if we compared the unconditional performance of method B with the conditional performance of method A (which corresponds to discarding data sets for the failing methods), this would reflect poorly on method B, even though both perform equally well for the data sets without failure.
+Clearly, we want to compare the methods’ unconditional performances.
+However, in both Tables 3(b) (a) and 3(b) (b), the relative unconditional performance of methods cannot be compared because the unconditional performance of method A is undefined.
+In contrast, the relative conditional performance can be quantified and compared (“discarding data sets associated with failure for all methods”): in Table 3(b) (a), the two methods have the same bias conditional on non-failure of method A; in Table 3(b) (b), the two methods have the same accuracy conditional on non-failure of method A.
+However, the conditional performance of method B should be of little interest.
+Why should a method’s performance depend on the behavior of another method regarding failure?
+In Table 3(b), method B’s performance improves when assessed based on only those data sets where A does not fail.
+This becomes even more complex as additional methods are included in the comparison, since discarding the corresponding data sets for all methods requires conditioning on the non-failure of each one.
+Alternatively, one could intuitively proceed to compare the unconditional performance of method B with the conditional performance of method A (“discard data sets for the failing methods only”).
+However, comparisons should be performed vertically, or we are not comparing like with like: In the context of Table 3(b), this would reflect poorly on method B, even though both perform equally well for the data sets without failure.
 This is all the more problematic since method B, in this case, is actually preferable to method A, as it outputs a result for all data sets!
-The methods’ conditional performance (“discard corresponding data sets for all methods”) does compare like-with-like.
-In conducting this comparison, method B’s conditional performance is in both cases (a) and (b) better than its unconditional performance.
-3.2 Caution with discarding data and imputation
-It is clear that what we would really like to quantify, but cannot, is relative unconditional performance.
-The superficial solution often proposed when looking at results such as those in Table 3(b) is to consider undefined values due to method failure as missing values and use missing data methods, such as imputation to infer the unconditional performance of method A.
-After all, results given by different methods tend to be positively correlated, so results given by method B could inform the imputation.
+Again, this becomes even more pronounced as additional methods are added to the comparison, since different methods typically fail on different datasets (see Section 3.1.1).
+2.2.2 Why imputation is inadequate
+Another superficial solution often proposed when looking at results such as those in Table 3(b) is to interpret method failure as causing missing values and use missing data methods, such as imputation, to infer the unconditional performance of method A.
+After all, results given by different methods tend to be positively correlated, so results given by method B could inform the imputation (analogously, it may be intuitive to impute with the results of method A on the remaining data sets as results tend to be correlated).
 We view it as mistaken to regard method failure as a form of missing data and have been careful to refer to “undefined” rather than “missing” values.
-The missing data literature defines missing values as those that exist but whose values were not recorded (Carpenter et al., , 2023; Rubin, , 2004).
+The missing data literature defines missing values as those that exist and were intended to be recorded but for one reason or another are “not observed” (Rubin, , 2004; Carpenter et al., , 2023).
 Method failure is far closer to what is called a truncating event in the estimands literature (Kurland et al., , 2009; Kahan et al., , 2024), which denotes a value that does not exist.
 That is, when a method fails, the result we wished to record simply does not exist for that data set.
-For this reason, we should be cautious about applying missing data approaches to method failure.
-In particular, imputing undefined performance values or discarding data sets associated with method failure (either for all methods or the failing methods only) do not enable the comparison of relative unconditional performance.
-There are two exceptions to this.
-First is when a “failure” is recorded but a result could have been obtained; for example, if failure occurs due to excessive runtime but the definition of excessive in the comparison study can safely be judged as shorter than in practice.
-For cases like this, it may make sense to run the experiment on a randomly selected subset of all simulation repetitions/benchmarking data sets for the method affected by runtime issues (“failure by design”).
-As this can be seen as missingness mechanism “missing completely at random” (missingness is “unrelated to the observed and unobserved data” (Carpenter et al., , 2023)), discarding the data sets for which the method “fails” or imputing based on the performance results from other methods or data sets may be appropriate.
-To ensure missingness completely at random, it is important to select the subset of data independently of the data sets for which the method has particularly long runtimes.
-Researchers should therefore think about possible runtime problems and carry out the corresponding random selection of data sets before conducting the study (see Section 3.3.7).
-A second exception is when failure occurs for no reason, such that it is safe to compare the unconditional performance of method B with the conditional performance of method A.
-However, because method failure always happens for a reason – the interplay of a particular method with a particular data set – we do not believe this would ever occur.
-An often more suitable approach to handling method failure than discarding data sets or imputing, which we term the “fallback strategy”, involves specifying a backup method when a failure occurs.
-This strategy is described in more detail in Section 3.3.5.
-Table 4: Recommendations on handling method failure
-Caution with Explanation
-Imputation In most cases: method failure does not lead to “missing data”
-$\rightarrow$ Possible exception: Runtime issues
-Discarding data
-(for all methods or the
-failing methods only)
-Usually does not provide assessment and comparison of relative unconditional performance
-$\rightarrow$ Possible exception: Runtime issues
-Alternative recommendations Explanation
-Pinpoint interplay causing failure as the basis of any of the following considerations…
-Check for method misuse Is the failing method truly developed for given estimation or prediction task?
-Check if data matches scope Are targeted users likely to encounter this kind of data and failure?
-Be realistic when modifying method implementations Consider changes in implementations to solve method failure that are realistic for targeted users and research stage.
-Possibly combine with fallback strategies (see below).
-Incorporate fallback strategy Incorporate method pipelines “method A if it succeeds in producing an output, method B otherwise”.
-$\rightarrow$ Enables assessment of relative unconditional performance.
-Report transparently … - failure proportions, ideally as performance measure.
-- interplay causing failure with sufficient detail: “Method Z, with an overall failure proportion of $X\%$, fails when … and when … .” - chosen handling(s) and justify choices based on above considerations.
-Slimmed-down recommendations
-Analyze and report simultaneously:
-(i) results after discarding data sets for all methods.
-(ii) results after discarding data sets for the failing methods.
-(iii) proportions of method failure for all methods.
-$\rightarrow$ Report that this does not provide assessment of relative unconditional performance.
-3.3 Practical recommendations in the presence of method failure
-In the following, we present practical recommendations in the presence of method failure that may be more appropriate than imputing or discarding data sets.
-We advise implementing these recommendations as comprehensively as possible as they contribute to increased usability of comparison studies to users of the methods.
-However, we are aware that methodological researchers facing time constraints and page limits may be prevented from following them completely.
-In Section 3.4, we therefore discuss how a “slimmed-down” version of our recommendations can be applied nevertheless.
-3.3.1 Pinpoint interplay of method and data characteristic causing failure
-We recommend authors of comparison studies investigate and understand for which types of data sets and why a corresponding method fails, beyond the initially perceived symptom(s).
-Ideally, they should identify if particular (and if different) data characteristics cause the failures, leading to evaluations such as “method A fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes”.
-Having identified the dynamic between the given method and data that causes method failure, authors should base their further handling on the considerations and recommendations provided in the following Sections 3.3.23.3.7.
-3.3.2 Check for method misuse
-When encountering method failure, authors should first check whether the method was truly developed for the given estimation or prediction task, and the type of data (based on the investigations from the previous Section 3.3.1), and if it is used as intended by the developers.
-If not the case, what appears initially as method failure may actually be method misuse.
-Method misuse should not be attributed to the method but to the user, that is, the method should not be penalized for failing on a data set or a task it was not intended for.
-To avoid method misuse and any associated failure in the first place, authors of comparison studies should familiarize themselves thoroughly with the methods under investigation and read the corresponding user manuals carefully.
-It may also make sense to consult the developers of the individual methods (and possibly involve them in the study).
-Note that there may also be cases where a method is commonly misused by real-world users, especially when they are inexperienced.
-Depending on which kind of user the study aims at (e.g., inexperienced versus experienced), the occurrence of this kind of failure may be relevant to them and should therefore be included in the study.
-In these cases, the misuse should be clearly emphasized.
+Method failure should, therefore, not be viewed as obscuring the method’s results on a given data set.
+Instead, failure is the result we get!
+For this reason, we should be cautious about applying missing data approaches and particularly imputation to method failure.
+Note that this also includes measures of “weighted performance,” which weight the performance of a method according to the proportion of data sets with failure for that method.
+Using the correspondence between weighting and imputation approaches, this implies that weighting implicitly assigns some performance value in instances of method failure.
+The consequence of our dissuasion from “missing data-inspired” approaches is that our work focuses intensively on alternative ways to deal with method failure.
+Note that there is an exception to method failure not leading to “missing” data.
+This is when a “failure” is recorded but a result could have been obtained; for example, if failure occurs due to excessive runtime but the definition of excessive in the comparison study can safely be judged as shorter than in practice.
+We refer to this case in Section 3.1.6.
+3 Principles and recommendations for handling method failure
+The following section presents guidelines for handling method failure that are more appropriate than discarding data or imputation.
+Since, as previously mentioned, method failure can only be directly observed after the execution stage, the order of our recommendations aligns with the natural process of investigating these failures post-execution.
+However, we also highlight key considerations that can be addressed during the planning stage to help reduce both the potential for and unpredictability of failure in the first place.
+Using the terminology of Heinze et al., (2024), we distinguish between comparison studies in “early” and “late” research phases.
+Early phase studies investigate if a given method can be “used with caution” in settings similar or only slightly different from its original target setting, possibly refining or extending it while also making limited comparisons.
+Late phase studies, however, focus on realistic comparisons across diverse scenarios to guide real-world users.
+We make differentiated recommendations when necessary.
+3.1 Practical recommendations for the handling of method failure
+3.1.1 Consider method failure as the result of a complex interplay
+When looking beyond the manifestations of method failure summarized in Section 2.1.1, it becomes evident that these are just the tip of the iceberg.
+In the end, regardless of how method failure manifests, it is rooted in an individual and complex interplay of interconnected factors, that is
+certain data set characteristics cause the method to fail in its current implementation (and on the given hardware and software).
+Besides influencing runtime and memory consumption (Zimmermann, , 2020), data set characteristics contributing to method failure are highlighted by the observation that a method usually fails for specific data sets, not all of them.
+Certain data characteristics can complicate or completely prevent required calculations of a method within the maximum number of iterations (non-convergence).
+Recall the study of van Smeden et al., (2016) addressing perfect separation of the outcome variable by one or more predictor variables as a cause for non-convergence of logistic regression algorithms.
+Additionally, Fernández-Delgado et al., (2014), in their extensive comparison study of supervised classification methods, list “collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al., , 2014) as data set characteristics associated with method failure.
+Conversely, the association of method failure with the (implementation of) a given method becomes clear from the usual observation that not all methods fail for the same data sets.
+That is, methods often vary in their ability to handle specific data set characteristics.
+For instance, in our illustrative study in Section 4.1, some odds ratio estimation methods are equipped with a way around “zeros,” while others fail.
+In addition to the interplay of method and data set characteristics, however, software can also affect the (non-)failure of a method such that results are obtained in one software but not another.
+Consider, for instance, that Herrmann et al., (2021) reported that one of the methods under investigation was completely unusable due to an error in R that occurred using Windows as an operating system, while it did not occur under Ubuntu.
+The dependency on hardware, on the other hand, becomes clear in how working memory, besides being directly associated with possible memory issues, can also considerably affect runtimes.
+Ultimately, the core of method failure–stemming from the interplay of data set, method, software, and hardware–is considerably more informative for addressing method failure than the manifestation of failure itself.
+We, therefore, recommend that authors of comparison studies investigate and understand the interplay underlying failure for each instance after the execution phase.
+Ideally, they should identify if particular (and if different) data characteristics cause the failures, leading to evaluations such as “method A fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes.” As a preparatory measure in the planning stage, authors should check user manuals and help pages for any described problems related to certain data structures for all methods under investigation.
+3.1.2 Check for the correct use of the method
+When encountering method failure in the execution stage, authors should first check whether the method was truly developed for the given estimation or prediction task and the type of data (based on the investigations from the previous Section 3.1.1), and if it is used as intended by the developers.
+If this is not the case, a distinction should be made between two scenarios.
+First, a method in the study is used in a slightly different (but not contraindicated) context than originally intended by its developers.
+We call this “use beyond the original scope.” Including this method in the study may be justified when real users commonly use this method “beyond its original scope,” however, authors should make a corresponding remark in their manuscript.
+Second, if using the method in the given context is explicitly contraindicated by the method developers, including it in the study would correspond to method misuse.
+A method should not be penalized for failing on a data set or a task it was explicitly not intended for.
+Even if commonly (but probably unknowingly) “misused” by real users, this method should be excluded from the study, and the contraindication should be stated clearly in the manuscript.
+If none of the scenarios apply, authors should resort to Section 3.1.3.
+All assessments regarding the use of a given method should be based on the familiarization with all methods from the experiment, which should already take place in the planning stage.
+This underlines the importance of carefully reading corresponding user manuals and checking for publications from real-world applications in which these methods are commonly used.
+This aspect should also be checked when planning the data generation in simulation studies and selecting benchmark data for benchmark studies.
+When in doubt, authors may also consult the developers of the individual methods (and possibly involve them in the study).
 3.1.3 Check if data matches scope of study
-If method failure remains after solving or ruling out method misuse, authors should assess if real-world users are likely to encounter such data (and therefore, this method failure) in practice.
-These considerations must include the scope of the study and are of course not purely objective.
-Importantly, they should be reported in the study, since they carry valuable information and affect the interpretation of the study’s results.
-Consider the case of method failure caused by the method’s inability to make required calculations due to an “extreme” structure of the given simulated data set, for example highly imbalanced classes in the outcome or predictor variable(s), as in White et al., (2024).
-Depending on the scope of the study, these extreme data sets may be of interest or not.
-If they are deemed unrealistic for the considered scope and thus not relevant to the readers of the study, authors may consider making sensible parameter changes in the DGM to generate data that is more realistic given the aim of their study, in line with White et al., (2024).
-Similarly, “too extreme” data may also affect real data based studies.
+If method failure persists after ruling out an incorrect use in the execution stage, authors should assess whether real-world users are likely to encounter such data (and, therefore, this method failure).
+This assessment is shaped by the study’s scope and is naturally not purely objective.
+Importantly, associated considerations should be reported in the study since they carry valuable information and affect the interpretation of the study’s results.
+Consider the case of method failure caused by the method’s inability to make required calculations due to an “extreme” structure of the given simulated data set.
+For instance, White et al., (2024) conduct an illustrative simulation study that examines methods for handling missing values in confounders in epidemiological data.
+They observe failure of an imputation method, and investigations of the corresponding data reveal that it occurs when outcome or exposure are highly imbalanced for individuals with an observed confounding variable.
+From this information, they conclude that “the data-generating mechanism is too extreme and should be changed to generate more outcome events” (White et al., , 2024).
+If, given the scope of the study, certain data set characteristics are deemed unrealistic for the targeted readers, authors may consider making sensible parameter changes in the DGM to generate data that is more realistic, in line with White et al., (2024).
+Similarly, “too extreme” data may also affect real data-based studies.
 If one of the (real) benchmarking data sets is considered to be irrelevant to the scope of the study after careful examination, authors may consider removing the data set from the study completely.
 See Section 4.2 for an illustration.
-However, if the data sets causing failure are deemed relevant to the study’s scope, the researchers conducting the comparison study should—more appropriate than discarding or imputing in most cases—proceed by considering how real-world users are likely to behave in this situation.
+However, if the data sets causing failure are deemed relevant to the (simulation or benchmark) study’s scope, the researchers conducting the comparison study should—more appropriate than discarding or imputing in most cases—proceed by considering how real-world users are likely to behave in this situation.
 In particular, they may modify certain parameters of the concerned method(s) (see Section 3.1.4) or resort to a fallback (see Section 3.1.5).
+To be able to assess whether data structures associated with failure are too extreme, certain considerations should already be made in the planning stage.
+They should be aware of the scope and goal of their study: is it the goal to examine the general behavior of the method(s) (i.e., earlier stages of research) or to provide concrete recommendations for real-world practitioners (later stages of research)?
+Is the focus on testing the methods under regular or more extreme circumstances?
+If aiming to provide recommendations to real-world users, what degree of extremity in the data remains realistic for them?
 3.1.4 Be realistic when modifying method implementations
-As already touched upon, in some instances of method failure, a way of mimicking the behavior of real-world users may be to resort to an alternative implementation of the affected method (see Masaoud and Stryhn, (2010) for an example).
-This way, method failure may be resolved completely, allowing for the assessment of unconditional (absolute and relative) performance of the methods under investigation.
-For instance, one could modify a parameter of the method, such as increasing the maximum number of iterations in the iterative maximum likelihood process or changing the internally employed optimization algorithm.
-However, authors of comparison studies should not tweak the implementation of the method thoughtlessly.
-Instead, they should always consider if the applied change in the parameter setting fits the given research stage and corresponding scope of the study.
-For instance, if the study goal is to understand the methods’ behaviors quite generally (and in earlier phases of research according to the terminology of Heinze et al., (2024)), authors may modify a greater variety of parameters to see when the method is successful and when it fails.
-However, if the study takes place in later phases and aims to guide inexperienced users, the applied parameter changes should be more restricted to what a naive user of the method would be able and likely to do.
-In any case, the performance assessment of the given method is only meaningful if the authors stress and justify their parameter choices when reporting the study.
-Furthermore, it is important to make sure that the method comparison is still fair, meaning that the effort spent on each method under consideration should not differ significantly between the methods, especially in late phase studies, where neutrality is of crucial importance.
-Finally, one should be careful about increasing the runtimes of the affected method(s) when running a comparison study.
-Especially when the number of iterations is raised, runtime or memory consumption can increase notably and might result in runtime or memory issues so that method failure is, after all, not solved.
-If, after these considerations, a change in the implementation of the given method is deemed reasonable, any deviation from the default parameter setting of a given method should be reported and justified transparently.
-Note that, as will be discussed in Section 3.3.5, changes in a method’s implementation could also be applied as a fallback strategy.
+In earlier stages of research, it may make sense to modify a method’s implementation entirely (i.e., for all simulation repetitions/benchmark data sets) when encountering method failure in the execution stage.
+For instance, authors may conclude after execution that the default number of maximum likelihood iterations is generally too low, or that the given optimization algorithm may fail frequently, and that an alternative optimizer is more reliable.
+This way, authors may modify a greater variety of parameters to see when the method is successful and when it fails.
+Although this early-stage research is not yet directly aimed at real-world users, authors should remember that it will be used by them later.
+Therefore, certain practical limitations on parameter modifications should apply to ensure that the method remains feasible for future users.
+Also, authors should be careful about increasing the runtimes of affected method(s).
+Especially raising the number of iterations in optimization algorithms can increase runtime or memory consumption, potentially shifting method failure rather than solving it.
+A special case of modifying a method’s implementation often only arises after the execution stage of the study: investigations following the failure of a method reveal code bugs in the implementation of a method.
+Especially when their programming experience is high, authors may fix this bug and notify the method developers.
+Ultimately, the performance assessment of the given method is only meaningful if the authors stress and justify their parameter choices when reporting the study.
 3.1.5 Consider a fallback strategy
-When contemplating ways to mimic real-world users when encountering method failure in later research stages, Morris et al., (2019) argue that they would usually not give up but naturally resort to an alternative method, a fallback, that successfully provides an output.
-Therefore, it can make sense to transfer this fallback strategy into the context of methodological comparison studies and evaluate the performance of method pipelines of the type
-“use method A if it successfully produces an output, use method B otherwise”.
-On the one hand, this approach incorporates what data analysts really do in practice, and on the other, it indirectly yields automatic handling of undefined performance values in case they affect method A but not method B.
-If chosen carefully and in a way that there is a suitable fallback for each instance of method failure, fallback strategies allow for the assessment of (absolute and relative) unconditional performance of the methods under investigation.
-Besides realistically reflecting the behavior of real-world users, this approach results in a performance assessment that depends only on the definition of the pipeline.
+When a comparison study aims to provide guidance to real-world users (i.e., in later phases of research), it makes sense to mimic their behavior when encountering method failure in the execution stage.
+How would they proceed when their method of choice failed on their data set?
+Morris et al., (2019) argue that they would usually not give up but naturally resort to an alternative method, a fallback, that successfully provides an output.
+Therefore, it can make sense to transfer this fallback strategy into the context of method comparison studies and evaluate the performance of method pipelines of the type
+“use method A if it successfully produces an output, use method B otherwise.”
+Incorporating fallbacks indirectly yields automatic handling of method failure in case they affect method A but not method B, allowing for the comparison of (absolute and relative) unconditional performance when chosen carefully and in a way that there is a suitable fallback for each instance of method failure.
 Note that when using fallback strategies, we no longer evaluate the “pure” performance of method A.
-However, it is clear from Section 3.1 that this in an unachievable goal in the first place.
-The concept of fallback strategies is already intuitively considered in some existing comparison studies.
-Voeltz, (2024), for instance, conclude from their study that “the approach by Frömke et al., (2022) could be used as a fallback strategy in case of non-convergence [of the overall best-performing method] because it always yields results”, where the approach by Frömke et al., (2022) is a non-iterative procedure, therefore never affected by non-convergence by design.
-Going one step further than Voeltz, (2024), one could provide a performance assessment for the resulting pipeline.
-A special case of the fallback strategy has indirectly been addressed in Section 3.3.4: it may sometimes make sense to define an alternative implementation of the given method (e.g, using a different optimizer or an increased number of iterations) as a fallback instead of a completely different method (as in Crowther et al., (2014) who changed the number of quadrature points).
-The corresponding pipeline would then consist of the default implementation of that method for those data sets where it successfully produces an output and the alternative implementation for the remaining data sets.
-In practice, the choice of a suitable fallback depends on the research phase.
-For instance, in early phases, the fallback may be constituted of an alternative implementation (e.g., alternative parameter choice) of the original method.
-Also, authors should strongly base the selection of fallbacks on what can be considered realistic given the expertise in statistics and programming of the targeted users.
-For instance, in a study targeting data analysts with limited statistical background, the fallback strategy may be an alternative method with default parameters rather than the same method with complex parameter settings.
-Since the choice of suitable fallbacks involves some subjectivity, authors should justify their choice(s) transparently.
-If failures are expected at the planning stage, i.e., before conducting the study, one may consider the whole pipeline as a competing method to be assessed in the comparison study.
-Another approach would be to design a pipeline after seeing the results of the comparison study, as did by Voeltz, (2024).
-In this case, the definition of the pipeline can be viewed as a result of the comparison study rather than a method that has to be evaluated in the first place.
-The definition of the fallback strategy, however, may not always be as straightforward as in Voeltz, (2024).
-It may for example make sense to consider complex rules involving a second fallback in case the first one fails.
-3.3.6 Transparently report failure proportions, underlying interplays, and applied approaches
-Following all recommendations above, authors should transparently report (i) any occurrences and proportions of method failure, (ii) the underlying interplay(s) causing failure (in different instances), and (iii) the corresponding approaches following the in-depth investigations of method failure.
-Report any occurrences and proportions of method failure  This can serve as an important criterion for real-world users selecting a suitable method.
+However, it is clear from Section 2.2.1 that this is unattainable in the first place.
+The choice of a suitable fallback, again, depends on the scope of the study and the targeted users.
+Should it serve as a guide for experienced or rather inexperienced users?
+When targeting inexperienced real-world users, authors may choose a completely different method as fallback.
+For instance, Zapf et al., (2024) suggest from the results of their study that “the approach by Frömke et al (2022) could be used as a fallback strategy in case of non-convergence [of the overall best-performing method] because it always yields results” (Zapf et al., , 2024), where the approach by Frömke et al., (2022) is a non-iterative procedure, therefore never affected by non-convergence by design.
+When addressing experienced users, an alternative implementation of the given method (in variation of Section 3.1.4) may make a suitable fallback.
+Many considerations and decisions regarding fallback strategies should already be made in the planning stage of a study.
+This includes, first and foremost, if using fallbacks is suitable to the given study, and based on this decision and the scope of the study, whether a completely new method or a modified implementation should be used as fallback.
+Authors should further be aware that the neutrality principle of late phase comparison studies should always apply, that is, authors should avoid putting considerably more effort into only a few methods when choosing fallbacks for them.
+This becomes clear, for instance, when modifying the implementation of a method, as, e.g., parameter modifications can strongly influence performance (Weber et al., , 2019).
+In the same context, it is important to avoid selective reporting as it can lead to over-optimistic results; that is, authors should disclose all tested pipelines, not just the best ones.
+It is therefore advisable to reduce the data-driven construction and the complexity of pipelines, which can be achieved by the following considerations.
+First, authors should decide on a maximum number of fallbacks that can be considered to realistically reflect the behavior of real users (as they are unlikely to try out an indefinite number of methods).
+Suitable combinations of methods and fallbacks may be preliminarily constructed based on the similarity of ease of use (e.g., a user is unlikely to resort to a fallback that is considerably more difficult to apply than the original method).
+Alternatively, inspired by Zapf et al., (2024), it may make sense to define a default fallback to all instances of method failure if a method has proven to be particularly computationally robust in previous comparison studies.
+Finally, authors should be aware that instances of all fallbacks failing on a given data set cannot always be avoided, even when using a computationally robust “default” fallback.
+Many fallbacks (and generally many methods) failing on the same data set may indicate that it is particularly “challenging.” Authors should then investigate the data set closely to check whether it suits the scope of the study (see Section 3.1.3) and consider removing it if it is deemed unrealistic for real users.
+Otherwise, reporting the corresponding characteristics of the data set alongside the methods that successfully produce an output may provide valuable insights to real users.
+3.1.6 Special case: Runtime issues
+In contrast to non-convergence and memory issues, method failure is actively declared by the authors when runtime issues occur in the execution stage.
+That is, running the methods under investigation on hundreds or thousands of data sets requires time restrictions, which are set by the authors arbitrarily, and without which the method would likely produce a valid output.
+On the other hand, it can sometimes be safely said that real (targeted) users, running the method on a single data set only, are likely to obtain an output.
+In these cases, when method failure leads to missing rather than undefined performance values, it may make sense to run the experiment on a randomly selected subset of all simulation repetitions/benchmarking data sets for the method affected by runtime issues (“failure by design”).
+To ensure that this “failure” is independent of data characteristics, it is important to select the subset of data independently of the data sets for which the method has particularly long runtimes.
+Researchers should therefore think about possible runtime problems and carry out the corresponding random selection of data sets in already the planning stage.
+3.1.7 Report failure proportions, underlying interplays, applied approaches, and share code
+Authors of comparison studies should report any occurrences and proportions of method failure.
+This can serve as an important criterion for real-world users selecting a suitable method.
 Even the “best performing” method is of limited use if the amount of settings in which it can be applied without frequent failures is limited.
-Ideally, the proportion of method failure should be included as an additional performance measure for all considered methods (see Morris et al., (2019)); see Voeltz, (2024) for an example.
-If method failure is predictable for certain settings and all affected methods fail for the same data sets (see, e.g., the example from Section 4.1), it may make sense to include the binary performance measure of the type “can handle setting: yes.
-vs.
-no” and report it alongside the corresponding number of repetitions/benchmark data sets associated with the failure.
-Report interplay of method and data causing failure  Adding to the previous paragraph, a corresponding reporting for each method can look like “method A, with a failure proportion of $X\%$, fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes”.
-This enables a more comprehensive overview of the applicability of the methods compared to reporting the corresponding failure proportions only.
-Readers can learn in which settings a method is applicable, and no less importantly, in which it is not.
+Ideally, the proportion of method failure should be included as an additional performance measure for all considered methods (see Morris et al., (2019)); see Zapf et al., (2024) for an example.
+Additionally, the interplay of method and data causing failure should be reported.
+Adding to the previous paragraph, a corresponding reporting for each method can look like “method A, with a failure proportion of $X\%$, fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes.” In addition to reporting the corresponding failure proportions, readers can learn in which settings a method is applicable, and no less importantly, in which it is not.
 Additionally, these insights serve as justification of the chosen handlings to method failure.
-An example of a corresponding reporting can be found in Hornung et al., (2024) who provide detailed descriptions in the supplement of their paper.
-Report applied handlings  Any actions resulting from the above-described in-depth investigations and recommendations regarding method failure should be reported with sufficient detail.
-This includes the removal of any misused methods, the removal of data (or modification of DGMs) that do not match the scope of the study, any changes in implementations, and the incorporations of fallback strategies.
+An example of a corresponding reporting can be found in Hornung et al., (2024), who provide detailed descriptions in the supplement of their paper.
+Authors of comparison studies should additionally report their applied handlings.
+Any actions resulting from the above-described in-depth investigations and recommendations regarding method failure should be reported with sufficient detail.
+This includes removing any misused methods, removing data (or modification of DGMs) that do not match the scope of the study, any changes in implementations, and incorporating fallback strategies.
 Additionally, authors should justify their choices given the scope of the study (i.e., which users it is aimed at), and the research stage in which it takes place.
-3.3.7 Finally: think ahead, pre-register but remain flexible
+The complexity of the occurrence and handling of method failure, which becomes clear from the previous sections, emphasizes all the more the value of sharing code, data, and intermediate results.
+Allowing authors of other, “new” comparison studies in-depth investigations to the specific instances of method failure in the given study may help them to a greater understanding of this topic–and enhance fruitful discussions on suitable handlings.
+3.1.8 Finally: think ahead and pre-register, but remain flexible
 The previous examples and those that will be outlined in Section 4 make clear that method failure and its cause(s), related to complex dynamics between factors such as method and data set, are often difficult to anticipate.
-While pre-specifying the handling of undefined values has the advantage of reducing analytical uncertainty and its dangers in terms of selective reporting (Siepe et al., , 2024), pre-specifying a standard procedure prior to conducting the study may sometimes be delicate.
-Our recommendation is, as far as possible, to consider in advance which combinations of data and methods could lead to failure.
-For instance, carefully reading user manuals and help pages of all methods before conducting the study might reveal specific requirements of a method on the data.
-This may make some instances of method failure less unexpected.
-More generally, making method failures less unexpected can be achieved by conducting a pilot experiment.
-Pilot experiments test whether an experimental design is appropriate for the specific context in which it is applied, and are recommended in several textbooks on Design and Analysis of Experiments (Cohen, , 1995; Dean et al., , 2017; Montgomery, , 2020).
-For example, in the context of real data comparison studies, a pilot experiment might use a fraction of the available data sets, say $20\%$, to test whether the specified method implementations produce reasonable results.
-The remaining $80\%$ of the data sets are then used for the main experiment, which may be refined based on what is learned in the pilot.
-A pilot study can also help authors anticipate possible runtime issues for which they can then pre-specify the random subset of data sets on which the experiment for the affected method will be restricted.
-Of course, it is important to clearly report the results of the pilot study too.
-After implementing these strategies, pre-specifying a procedure for handling method failure may be less hazardous.
-However, even despite pre-registration, researchers should be aware that unforeseen issues may appear and require adaptations of the pre-specified strategy.
-See Lakens, (2024) for a discussion of the consequences of deviations from pre-registration on the severity and validity of inferences.
-3.4 Slimmed-down recommendations when time is short
+While pre-specifying the handling of undefined values has the advantage of reducing analytical uncertainty and its dangers in terms of selective reporting (Siepe et al., , 2024), pre-specifying a standard procedure before conducting the study may sometimes be delicate.
+Making method failure less unexpected (and pre-specifying a procedure for handling method failure less hazardous) can be achieved through pilot experiments.
+They test whether an experimental design is appropriate for the specific context in which it is applied, and are recommended in several textbooks on Design and Analysis of Experiments (Cohen, , 1995; Dean et al., , 2017; Montgomery, , 2020).
+For example, a pilot experiment might use a fraction of the available data sets (e.g., $20\%$) in the comparison study to test whether the specified method implementations produce reasonable results.
+The remaining $80\%$ is used for the main experiment, which may be refined based on what is learned in the pilot.
+Pilot studies can also help authors anticipate possible method failures for which they can then pre-specify adequate handlings.
+Note that it is important to clearly report the results of the pilot study.
+Despite pre-registration, researchers should be aware that unforeseen issues may appear and require adaptations of the pre-specified strategy (see Lakens, (2024) for a discussion of the consequences of deviations from pre-registration on the severity and validity of inferences).
+As a result, error-handling mechanisms should be included in the code right from the beginning, including the storage of error messages (White et al., , 2024).
+3.2 Slimmed-down recommendations for the busy researcher (looking to do just enough)
 For some authors, there may be time constraints that do not allow for any (in-depth) implementations of our recommendations from the previous Section 3.1, however, authors nevertheless wish to aggregate results over data sets despite the existence of undefined values caused by method failure.
-In these cases, where it is difficult to follow the entirety of our recommendations, we recommend the following slimmed-down analysis and reporting strategy.
+While we recommend implementing our in-depth recommendations as comprehensively as possible, we suggest the following slimmed-down analysis and reporting strategy when time and resources are limited.
 Authors may present performance results after (i) discarding the corresponding data sets for the failing methods, (ii) discarding these data sets for all methods, and (iii) report the failure proportions for all methods, ideally as an additional performance measure.
 This three-fold analysis and reporting enables more differentiated assessments of the methods’ performances and avoids the preference of methods based solely on how data sets associated with method failure are discarded.
-In particular, it provides insights into how the methods successfully producing an output perform on the data sets for which others fail (and how this can influence the method comparison).
-However, following Section 3.1, authors should be aware and clearly state in their manuscript that any form of aggregating performance results despite existing undefined values does not allow an assessment of unconditional performance of all methods under investigation.
+It also reveals how successful methods perform on data sets for which others fail (and how this influences the method comparison).
+When the graphical comparison involves boxplots of performance values, authors could indicate method failure (and the corresponding frequency) as scatters of contrasting color above or below the box.
+For instance, for bias, they could be indicated above the plot (as a value higher than all biases observed in the study), whereas for accuracy, it could be indicated below the plot.
+However, following Section 2.2.1, authors should be aware and clearly state in their manuscript that any form of aggregating performance results despite existing undefined values does not allow an assessment of unconditional performance of all methods under investigation.
 4 Illustrations
-In the following, we illustrate the guidelines presented in Section 3 through two illustrative comparison studies featuring method failure: a statistical simulation study and a benchmark study at the interface between statistics and predictive modeling.
-For both examples, we first demonstrate how the ad hoc approaches discussed in Section 3.2, such as discarding data sets associated with failure, can distort the results and even lead to misleading method recommendations.
-Then, we illustrate the implementation of the recommendations from Section 3.3.
-Thereby, our focus is exclusively on the handling of undefined performance values and its consequences on the method comparison, rather than the method comparison itself.
-This has three consequences.
-First, both illustrations represent only a small excerpt (e.g., a few or a single scenario) of what would be a typical publishable study.
-Second, we do not give a comprehensive overview of all aspects of the studies but focus on those that are relevant to the issue of undefined performance values.
-Third, our goal is not to provide insights into the performance of the compared methods (larger dedicated studies would be necessary for this purpose).
+In the following, we illustrate the guidelines presented in Section 3 through two fictive comparison studies featuring method failure: a statistical simulation study and a benchmark study at the interface between statistics and predictive modeling.
+Note that our focus is exclusively on the handling of undefined performance values and its consequences on the method comparison rather than the method comparison itself.
+This means that both illustrations represent only a small excerpt (e.g., a few or a single scenario) of what would be a typical publishable study, exclusively focusing on those aspects of the studies that are relevant to the issue of method failure.
+Importantly, our goal is not to provide insights into the actual performance of the compared methods.
 4.1 Example 1: Comparing methods for the estimation of odds ratios
 4.1.1 General
-Our first example, a statistical simulation study, deals with the comparison of odds ratio estimation methods.
-Odds ratios (“ORs”) quantify the association between an exposure and an outcome and are typically used in (retrospective) case-control studies.
-Here, we focus on the case of a binary exposure $X\in\left\{0,1\right\}$ and binary outcome $Y\in\left\{0,1\right\}$.
-While a straightforward way of estimating the OR is to calculate $\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$ (where $n_{ij}$ is the observed number of subjects with $X=i$ and $Y=j$), other methods employ more complex modeling strategies.
-Note that ORs are bounded in $(0,\infty)$, where values $>1$ imply a positive association between exposure and outcome.
-With the true OR value being generally unknown in real data sets, OR estimation methods are typically compared using simulation studies under different scenarios, e.g., different sample sizes, underlying true OR values, and probabilities of exposure $p_{x}$.
-Especially when the sample size is low, the underlying true OR is far from $1$, or $p_{x}$ is close to 00 or $1$, some simulated data sets may not contain any subjects for one or more of the four outcome-exposure combinations.
-This results in $n_{ij}=0$ for the corresponding combination $i,j\in\{0,1\}$ for this data set.
-These zeros are called “sampling zeros” and cause computational issues (e.g., division by 00) for many of the available OR estimation methods.
-They lead to either returning an error or, depending on the position of the sampling zero in the contingency table, non-meaningful OR estimations of 00 or $\infty$”, resulting in undefined performance values in the corresponding data sets.
+This fictive statistical simulation study deals with the comparison of odds ratio estimation methods.
+Outcome and exposure are binary ($X,Y\in\left\{0,1\right\}$) and the observed number of subjects with $X=i$ and $Y=j$ is denoted as $n_{ij}$.
+The scenarios are defined based on different sample sizes $n.obs$, underlying true OR values, and probabilities of exposure $p_{x}$.
+Especially when the sample size is low, the underlying true OR is far from $1$, or $p_{x}$ is close to 00 or $1$, some simulated data sets may not contain any subjects for one or more of the four outcome-exposure combinations ($n_{ij}=0$).
+These “sampling zeros” cause calculation issues (e.g., division by 00) for many OR estimation methods, leading to either returning an error or, depending on the position of the sampling zero in the contingency table, non-meaningful OR estimations of 00 or $\infty$” for the corresponding data sets.
 In practice, users often apply the “Haldane-Anscombe correction” (Haldane, , 1956; Anscombe, , 1956) to handle sampling zeros in the OR estimation.
@@
 4.1.2 Design of the fictive study
-In our fictive simulation study, we vary the underlying true OR (OR $\in\{2,3,4,5\}$) and exposure probability $p_{x}$ ($p_{x}\in\{0.25,0.5\})$ while keeping the sample size at the fixed value of $n.obs=50$.
+We vary the underlying true OR (OR $\in\{2,3,4,5\}$) and exposure probability $p_{x}$ ($p_{x}\in\{0.25,0.5\})$ while keeping the sample size fixed at $n.obs=50$.
 This results in eight distinct simulation scenarios (see Table 4 for an overview).
-For each scenario, we simulate 100 000100000100\,000100 000 outcome-exposure data sets (where the exposure variable $X$ is generated using a Bernoulli distribution with probability $P(X=1)=p_{x}$ and outcome $Y$ is generated using the logistic function with the linear effect of exposure $X$ specified as $\beta_{X}=\log(\text{OR})$).
-Five OR estimation methods are applied to the simulated data to obtain a point estimate for the true OR, namely “Fisher”, “Midp”, “Small” (all three implemented in R package epitools (Aragon et al., , 2020)), “Woolf” (Woolf et al., , 1955) (implemented in R package pairwiseCI (Schaarschmidt and Gerhard, , 2019)), and “Manual”, which corresponds to the manual computation $\widehat{OR}=\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$.
-Note that we omit method “Wald” from the epitools package since its estimation output is identical to the manual computation (i.e., method “Manual”).
+For each scenario, we simulate 100 000100000100\,000100 000 outcome-exposure data sets.
+Five OR estimation methods are applied to the simulated data to obtain a point estimate for the true OR.
+“Manual” corresponds to the manual and straightforward computation $\widehat{OR}=\frac{n_{11}\cdot n_{00}}{n_{01}\cdot n_{10}}$.
+The remaining four methods, namely “Fisher,” “Midp,” “Small” (all three implemented in R package epitools (Aragon et al., , 2020)), and “Woolf” (Woolf et al., , 1955) (implemented in R package pairwiseCI (Schaarschmidt and Gerhard, , 2019)) employ more complex modeling strategies.
 From the five estimators under investigation, only Small and Woolf can handle sampling zeros internally.
-For each OR estimator and simulation scenario, bias on the $\log$-scale is computed as
-[…]
-and the MSE on the $\log$-scale is obtained as
-[…]
-Note that the $\log$-transformation ensures symmetry around the value of “no association”. For both performance measures, values closer to 00 indicate better performance (Morris et al., , 2019).
-4.1.3 Occurrence and handling of method failure
+For each simulation scenario, the estimators are compared regarding bias on the $\log$-scale ($\log$-transformation ensuring symmetry around the value of “no association”).
+Our analysis consists of two parts.
+First, we compare the performance of the OR estimation methods when discarding data sets with sampling zeros for all methods versus only for the failing methods (“ad hoc approaches”; Section 4.1.3).
+Though users often apply the Haldane-Anscombe correction when encountering sampling zeros, this first part illustrates how the preference of some methods may mainly be driven by the way data sets with sampling zeros are discarded.
+Second, we repeat the analysis implementing the principle of fallback strategies suggested in Section 3.1.5.
+This includes important considerations that must be made when choosing suitable fallbacks.
+For instance, for the manual estimation of the OR (method “Manual”), only the Haldane-Anscombe correction is applied as a fallback, assuming that a user applying the manual computation in the first place, which is very easy to implement, is unlikely to resort to a method that requires notably more effort.
+On the other hand, for methods Midp and Fisher for data sets with sampling zeros, we (i) either apply the Haldane-Anscombe correction, (ii) use method Small, or (iii) use method Woolf, assuming that corresponding users are more likely to employ fallbacks that are slightly more difficult to apply.
+Together with Woolf and Small, this leads to nine “methods” (“pure” methods or pipelines) being compared in total.
+The results are presented in Section 4.1.4.
+4.1.3 Occurrence of method failure and ad hoc handlings
 Table 4 displays the proportion of the 100 000100000100\,000100 000 simulated data sets containing at least one sampling zero across the eight simulation scenarios.
@@
 8 50 5 $0.5$ 1.27%
-Our analysis consists of two parts.
-First, we compare the performance of the OR estimation methods regarding performance measures bias and MSE when ignoring data sets with sampling zeros for all methods versus ignoring data sets only for the failing methods (“ad hoc approaches”).
-While we are aware that users often apply the Haldane-Anscombe correction when encountering sampling zeros, this first part illustrates how the preference of some methods may mainly be driven by the way data sets with sampling zeros are discarded.
-Second, we repeat the analysis implementing the principle of fallback strategies suggested in Section 3.3.5.
-As fallbacks for methods Midp and Fisher for data sets with sampling zeros, we (i) either apply the Haldane-Anscombe correction, (ii) use method Small, or (iii) use method Woolf.
-For the manual estimation of the OR (method “Manual”), we only apply the Haldane-Anscombe correction as a fallback, assuming that a user applying manual computation in the first place, which is very easy to implement, is unlikely to resort to a method that requires notably more effort.
-Together with Woolf and Small, nine “methods” (“pure” methods or pipelines) are compared in total and the results are presented in Section 4.1.4.
-Note that since we focus on the effect of handling method failure (instead of the actual results of the comparison), and this effect becomes clearer through examining the individual scenarios, we do not aggregate the performances across the eight scenarios in either of the two analysis parts.
-Figure 1: Performance ranks of the OR estimation methods across the eight simulation scenarios, derived from the empirical bias in subfigure (A) and from the MSE in subfigure (B).
-In both subfigures, the ranks are compared between discarding data sets yielding failure for the failing methods only (denoted by “Failing” on the $x$-axis) versus for all methods (denoted by “All” on the $x$-axis).
-A rank of 1 means that the method has the lowest empirical bias (A) or MSE (B) in the given scenario.
-Note that Small and Woolf are the only methods in this selection that can deal with sampling zeros internally.
+Figure 1: Performance ranks of the OR estimation methods across the eight simulation scenarios based on empirical bias on the $\log$-scale, where a rank of 1 means that the method has the lowest empirical bias in the given scenario.
+(A) Ranks are compared between discarding data sets for the individual failing methods only (“Single” on the $x$-axis) versus for all methods (“All” on the $x$-axis).
+(B) The ranks are compared after applying fallback strategies to methods Midp, Fisher, and Manual.
+For Midp and Small, the fallbacks Small, Woolf, and the Haldane-Anscombe correction (“$+0.5$”) are used.
+For Manual, only the Haldane-Anscombe correction is used as fallback.
+Small and Woolf can deal with sampling zeros internally.
 Abbreviation of methods: “Fis”: Fisher; “Mid”: Midp“; “Sma”: Small; “Woo”: Woolf; “Man”: Manual (i.e., manual estimation of the OR).
-Ad hoc approaches  For the first analysis, Figure 1 shows that the way of discarding data sets with sampling zeros can substantially influence the results of the method comparison.
-This particularly affects estimator Woolf, one of the two methods that can handle sampling zeros internally.
-In terms of bias (Figure 1 (A)), its performance rank deteriorates in most scenarios and by up to three positions when discarding data sets with sampling zeros for all methods (including Woolf and Small).
-The reason for this is that across the simulated data sets without sampling zeros, Woolf slightly underestimates the true OR, while for the few data sets with sampling zeros, Woolf notably overestimates it.
-This leads to top performance when including all data sets in its performance assessment (while discarding data sets with sampling zeros for the failing methods).
-Even more so, the manner of discarding data sets with sampling zeros affects Woolf’s performance rank regarding MSE (Figure 1 (B)).
-In some scenarios, Woolf’s rank improves from bottom to top rank, indicating its poorer performance for data sets with sampling zeros.
-The performance rank of Small, the second method that can handle sampling zeros internally, is less affected by this for both bias and MSE.
-In particular, it scores last in terms of bias across all scenarios and for both manners of discarding (Figure 1 (A)).
-However, its performance rank changes by up to two positions for MSE (Figure 1 (B)).
-The performance ranks of the remaining methods are also affected regarding bias and MSE, where the manner of discarding data sets with sampling zeros affects their performance ranks by up to two positions.
-Note that the manner of discarding data sets has a larger impact on the performance comparison for scenarios with a higher proportion of sampling zeros.
-In particular, scenarios 5 and 6, with a maximum of $0.11\%$ of data sets containing sampling zeros, yield equal performance rankings across both handlings.
+(B): The combination of original method $m_{1}$ and fallback $m_{2}$ is marked by “$m_{1}/m_{2}$.”
+Ad hoc approaches  For the first analysis, Figure 1(A) shows that the way of discarding data sets with sampling zeros can substantially influence the results of the method comparison.
+This can lead to changes in the performance rank by up to three positions in some scenarios (see method Woolf) and particularly affect which method is chosen as the best.
+Note that a more detailed investigation of Woolf’s performance shows that for the data sets without sampling zeros, Woolf slightly underestimates the true OR, while for the few data sets with sampling zeros, Woolf notably overestimates it.
+This leads to top performance when all data sets are included in its performance assessment (while discarding data sets with sampling zeros for the failing methods).
+Generally, the manner of discarding data sets has a larger impact on the performance comparison for scenarios with a higher proportion of sampling zeros.
 However, differences in rankings within a given scenario are already observable with a proportion of $0.5\%$ of data sets with sampling zeros.
-Figure 2: Performance ranks of the OR estimation methods across the eight simulation scenarios when applying fallback strategies to methods Midp (“Mid”), Fisher (“Fis”), and Manual (“Man”).
-The rankings are derived from the empirical bias in subfigure (A) and from the MSE in subfigure (B).
-A rank of 1 means that the method has the lowest empirical bias (A) or MSE (B) in the given scenario.
-Note that for Midp and Small, the fallbacks Small (“Sma”), Woolf (“Woo”), and the Haldane-Anscombe correction (“$+0.5$”) are used.
-For Manual, only the Haldane-Anscombe correction is used as fallback.
-Small and Woolf are the only methods in this selection that can deal with sampling zeros internally, therefore not requiring any fallback.
-This amounts to nine “methods” in total.
-The combination of original method $m_{1}$ and fallback $m_{2}$ is marked by “$m_{1}/m_{2}$”.
-4.1.4 Recommended handling
-As expected, the relative performance between Woolf and Small remains the same between applying fallbacks for the other methods (Figure 2) versus discarding data sets with sampling zeros for the failing methods only (“failing” parts in Figure 1 (A) and (B)).
-This makes sense since neither method requires a fallback, resulting in unchanged bias and MSE.
-Midp’s overall good performance regarding bias and MSE is underpinned by using fallbacks Small and the Haldane-Anscombe correction.
-Here, it is debatable which fallback is preferable.
-Both fallbacks for Midp frequently switch ranking positions across the performance measures and simulation scenarios.
-However, fallback Small might be even easier to implement than the Haldane-Anscombe correction, requiring only a single parameter modification from Midp.
-Note the interesting observation that Small scores below the remaining methods regarding bias when applied alone, but performs well as a fallback for Midp.
+4.1.4 Recommended handling: using fallbacks
+Woolf’s overall good performance across all scenarios is underpinned when using fallbacks for all methods affected by method failure.
+Furthermore, the comparison between Woolf and Midp, another method performing well in most scenarios, becomes clearer through the use of fallbacks.
+While they often alternate in rank (and especially between rank 1 and 2) when removing data sets (Figure 1 (A)), their performance comparison is constant across all scenarios when a fallback is applied to Midp (Figure 1 (B)).
+Since this applies to both fallbacks, the Haldane-Anscombe correction and Small, it is debatable which fallback is preferable.
+Both fallbacks for Midp frequently switch ranking positions across the performance measures and simulation scenarios, however, fallback Small might be even easier to implement than the Haldane-Anscombe correction, requiring only a single parameter modification from Midp.
 4.2 Example 2: Comparing methods for computing confidence intervals for the generalization error
@@
 The “naive” method “$N$” for constructing the CIs ignores the dependence between the generalization performances in the resampling repetitions, while method “$C$” corrects for it.
-The performance of both methods is assessed in terms of coverage, which is the probability that a constructed CI includes the true value of the generalization performance.
-It should be close to its nominal value, e.g.
-$0.95$ if we consider $95\%$-CIs.
+The performance of both methods is assessed in terms of coverage.
+Note that a comprehensive benchmark study of methods for constructing CIs for the generalization error is performed by Schulz-Kümpel et al., (2024).
 4.2.2 Design of the fictive study
-Classification trees (Breiman et al., , 1984) are considered as models for predictive modeling.
-As benchmark data set, we use the data set provided by the SAPA project (Condon et al., , 2017), further processed according to Klau et al., (2023).
-The response variable is binary and five variables from the data set are used as predictors.
-The generalization performance is quantified using the Area Under the Curve (AUC), where a value of $0.5$ corresponds to a random prediction and a value of $1$ indicates perfect classification.
+Classification trees (Breiman et al., , 1984) are considered as models for predictive modeling and the data set provided by the SAPA project (Condon et al., , 2017), further processed according to Klau et al., (2023), is used as the benchmark data set.
+The response variable is binary.
+The generalization performance is quantified using the Area Under the Curve (AUC).
 The AUC is estimated using $15$ repetitions of repeated subsampling ($i=1,\ldots,15$) with a 4:1 ratio (i.e., the data is randomly split $15$ times into $80\%$ on which the model is trained and $20\%$ on which the AUC is obtained (Hastie et al., , 2009)).
@@
 $\Big{[}\ \overline{\widehat{AUC}}-t_{0.975,14}\cdot\sqrt{\Big{(}\frac{1}{15}+c\Big{)}\cdot S^{2}_{\widehat{AUC}_{i}}},\quad\overline{\widehat{AUC}}+t_{0.975,14}\cdot\sqrt{\Big{(}\frac{1}{15}+c\Big{)}\cdot S^{2}_{\widehat{AUC}_{i}}}\ \Big{]},$ (1)
-where $\overline{\widehat{AUC}}$ represents the mean AUC estimate over the $15$ subsampling repetitions and $S^{2}_{\widehat{AUC}_{i}}$ the corresponding sample variance, while $t_{0.975,14}$ is the $95\%$-quantile of the $t$-distribution with $15-1=14$ degrees of freedom.
-The fixed correction term $c>0$ incorporates the correlation structure between the subsampling repetitions (Nadeau and Bengio, , 2003).
-Method $N$, on the other hand, does not apply any correction, i.e.
-sets $c=0$ in the above formula.
-[…]
-The study is conducted using R package mlr3, method $N$ is implemented using function t.test() from package base and for method $C$, $\overline{\widehat{AUC}}$ and $S^{2}_{\widehat{AUC}_{i}}$ are obtained using functions mean() and sd() (from the same package).
+where $\overline{\widehat{AUC}}$ represents the mean AUC estimate and $S^{2}_{\widehat{AUC}_{i}}$ the sample variance over the $15$ subsampling repetitions.
+$t_{0.975,14}$ is the corresponding quantile of the $t$-distribution.
+A fixed correction term $c>0$ incorporates the correlation structure between the subsampling repetitions for method $C$ (see Nadeau and Bengio, (2003) for the construction of c).
+Method $N$, on the other hand, sets $c=0$ in the above formula.
+To estimate the coverages of methods $N$ and $C$, the following procedure is repeated for $1000$ iterations (see Figure 2).
+The SAPA data set is randomly split into a large test set $D_{test}$ ($80\%$ of observations) and a training set $D_{train}$ (20% of observations).
+The true AUC is approximated by applying the trained model to $D_{test}$.
+The corresponding estimated AUC, on the other hand, is obtained via repeated subsampling on $D_{train}$.
+Confidence intervals for the true AUC are constructed according to Equation  (1) with corresponding values of $c$ for methods $N$ and $C$.
+Finally, the estimated coverages of methods $N$ and $C$ are the proportions of the $1000$ iterations where the respective confidence intervals cover the approximated true AUC.
+The study is conducted in R using package mlr3, method $N$ is implemented using function t.test() from package base and for method $C$, $\overline{\widehat{AUC}}$ and $S^{2}_{\widehat{AUC}_{i}}$ are obtained using functions mean() and sd() (from the same package).
 Figure 2: Design of the fictive study to obtain the empirical coverage of methods $C$ and $N$. An approximation of the “true AUC” is obtained based on the very large $D_{test}$.
-4.2.3 Occurrence and handling of method failure
+4.2.3 Occurrence of method failure and ad hoc handlings
 For method $N$, the implementation of the above-described procedure returns the following error message for 300 of the $1000$ iterations in R:
@@
   • 3.
-    for method $N$, set “undefined” confidence intervals to “not covering the true AUC”.
+    for method $N$, set “undefined” confidence intervals to “not covering the true AUC.”
 Note that the intuition to approach 3, which can be viewed as a form of imputation, is that a CI that does not exist cannot cover any value.
@@
 The coverages resulting from the ad hoc approaches are displayed in Table 5 (a).
-While the coverage of method $C$ is constant across all three ad hoc approaches (coverages above the nominal level indicate conservativeness), method $N$, while too liberal across all approaches, is even more liberal using the ad hoc imputation approach.
+While the coverage of method $C$ is constant across all three ad hoc approaches, method $N$, while too liberal across all approaches, is even more liberal using the ad hoc imputation approach.
 4.2.4 Recommended handling
-We now closely inspect the modeling processes underlying the study.
-The goal is to find out the root cause of the error message and to adapt the handling of the undefined performance values accordingly.
-The errors in method $N$ occur when the estimated AUCs are equal across all $15$ subsampling repetitions, resulting in a sample variance $S^{2}_{\widehat{AUC}{i}}$ of 0.
+We now closely inspect the modeling processes underlying the study to find the root cause of the error message and to adapt the handling of method failure accordingly.
+Investigations show that the errors in method $N$ occur when the estimated AUCs are equal across all $15$ subsampling repetitions, resulting in a sample variance $S^{2}_{\widehat{AUC}{i}}$ of 0.
 This zero variance in the denominator of the $t$-statistic causes an error in function t.test().
-However, focusing on the confidence interval, we realize that from a theoretical point of view, a sample variance of 0 simply means that the CI is the zero-width interval $\left\{\overline{\widehat{AUC}}\right\}$.
+Focusing on the confidence interval, however, a sample variance of 0 simply implies a zero-width CI $\left\{\overline{\widehat{AUC}}\right\}$.
 Therefore, method $N$ not yielding a CI (and returning an error) is a technical artifact of its implementation.
 A straightforward solution would be to modify the implementation of method $N$, using mean() and sd() instead of t.test(), so that it computes the zero-width CI, which can enter the calculation of coverage as any other CI (see Table 5 (b) for the resulting coverage).
-However, taking the investigation another step further reveals that a sample variance $S^{2}_{\widehat{AUC}{i}}=0$ always occurs alongside all $15$ estimated (and the approximate true) AUC being equal to $0.5$.
+However, upon yet further investigations, it becomes clear that a sample variance $S^{2}_{\widehat{AUC}{i}}=0$ always coincides with all 15 estimated (and the approximate true) AUC being $0.5$.
 Ultimately, this worst possible performance of the classification tree (“random prediction”) occurs when the classification tree does not perform any splitting.
@@
 For example, the problem disappears if we include additional predictors: splitting then occurs, leading to an AUC higher than $0.5$.
-The adequate handling of method failure in this illustrative study, as described in Section 3, depends on the aim of the study.
-In particular, a prediction problem in which splitting often does not take place can be viewed as an extreme scenario.
-If the study’s scope includes such extreme scenarios, one could attempt to suitably modify parameters of the classification tree to favor splitting.
-If, however, the study aims to compare both methods in regular (i.e., less extreme) scenarios, one might consider additional predictors to favor splitting more strongly, resort to an alternative prediction model (e.g., logistic regression), an alternative measure of generalization performance (e.g., accuracy), or use a different benchmarking data set.
-Finally, note that this example illustrates that method failure may be complex and that the interplay causing it may not be immediately identifiable, even if the benchmark study consists of a single scenario only.
-In a real benchmark study which typically contains multiple scenarios, thorough investigations can become even more difficult.
+This illustrates well that the further the investigations are taken, the apparent adequate handling of method failure can change notably.
+In this case, based on the information gained, the adequate handling depends on the aim of the study and requires further considerations such as the following.
+A prediction problem where splitting often does not occur can be viewed as an edge case.
+If the study’s scope includes such edge cases, one could attempt to suitably modify the parameters of the classification tree to favor splitting.
+If, however, the study aims to compare both methods in regular scenarios, one might consider adding additional predictors to favor splitting more strongly, resort to an alternative prediction model (e.g., logistic regression), an alternative measure of generalization performance (e.g., accuracy), or use a different benchmarking data set.
 5 Discussion
-Method failure in comparison studies, often manifesting as issues regarding calculation (e.g., non-convergence), runtime, or memory issues, complicates the desired comparison of unconditional performance between the methods under investigation, i.e.
-across all simulated or benchmark data sets.
-Our informal literature review and illustrations both make clear that authors should not blindly rely on the initially perceived manifestation of failure but rather make extensive investigations of the sources of failure, which often guides the choice of the handling of the resulting undefined values.
+Method failure in comparison studies, often manifesting as issues regarding calculation (e.g., non-convergence), runtime, or memory issues, complicates the desired comparison of unconditional performance between the methods under investigation, i.e., across all simulated or benchmark data sets.
+Our informal literature review and illustrations both make clear that authors should not blindly rely on the initially perceived manifestation of failure but rather make extensive investigations of the underlying factors causing the failure, which often guides the choice of the handling of the resulting undefined values.
 Even if we argue for decisions on a case-by-case basis after careful inspection of the failures, we show and illustrate that some general principles hold in most situations.
-In particular, method failure usually leads to non-existing values instead of just “missing”, i.e., existent but unobserved values.
-This means imputation, a common but missing-data-inspired approach to handling method failure, is often inadequate.
-Another popular and seemingly logical approach, namely discarding data associated with failure (either for all or failing methods), almost never enables the comparison of unconditional performance.
+In particular, method failure usually leads to non-existing values instead of just “missing,” i.e., existent but unobserved values.
+This means imputation, a common but missing-data-inspired approach to handling method failure, is usually inadequate.
+The same applies to the popular and seemingly logical approach of discarding data associated with failure (either for all or failing methods), which almost never enables the comparison of unconditional performance.
 In this work, we provide alternative approaches that might be more appropriate to most occurrences of method failure.
-An approach that in our view deserves more attention than it receives currently in the literature consists in evaluating whole “pipelines” including one or several fallback strategies in case of method failure.
+An approach that in our view deserves more attention than it receives currently in the literature consists of evaluating whole “pipelines,” including one or several fallback strategies in case of method failure.
 This approach reflects the typical behavior of most users of methods in practice while enabling the comparison of unconditional performance if constructed carefully: if one method does not work, they resort to an alternative one.
@@
 When time resources and expertise are restricted, transparent reporting is all the more important.
+Furthermore, this underlines the necessity for openly sharing code, data, and ideally intermediate results, as this allows other researchers to explore the observed instances of method failure in greater detail.
 Also, our viewpoint on method failure as leading to “undefined” performance values does not always hold unrestrictedly, which can complicate the implementation of our recommendations.
@@
 One such example is near-separation (van Smeden et al., , 2016), which does not necessarily make itself obvious, hence the existence of dedicated methods to detect it (Mansournia et al., , 2017).
-Such grey-area failures add further complexities to benchmark and simulation studies: overall performance may depend on the specific criteria used to judge failure.
+Such gray-area failures add further complexities to benchmark and simulation studies: overall performance may depend on the specific criteria used to judge failure.
+Partial failure of a method is also possible.
+For example, in a simulation study, a method might return a valid point estimate but not its standard error, or a point estimate for one estimand but not another.
+Many of the principles discussed in this article apply in such situations, for example, the principle to avoid imputing estimates where a method has failed, and the notion of a pipeline (when the standard error estimator fails, revert to a backup estimator if available).
+There are, however, further subtleties.
+For example, it may be inadvisable for the fallback approach to keep the point estimate from method A and fallback on the standard error estimated from method B.
 To sum up, we believe there is not “one” right way to handle method failure in all situations.
@@
 They lead to helpful strategies for dealing with it, as outlined as part of our recommendations.
-However, we also believe that undefined performance values are only one of many issues related to comparison studies that deserve much more time, efforts and methodological strength towards more reliable evidence on the behavior of methods in the context of a “replication crisis in methodological research” (Boulesteix et al., , 2020).
+However, we also believe that method failure is only one of many issues related to comparison studies that deserve much more time, efforts, and methodological strength towards more reliable evidence on the behavior of methods in the context of a “replication crisis in methodological research” (Boulesteix et al., , 2020).
 We need more well-designed neutral comparison studies (Boulesteix et al., , 2017)—late phase studies according to the terminology of Heinze et al., (2024).
@@
 Acknowledgements
-The authors thank F.
-Julian D.
-Lange for useful literature input and Clemens Kreutz for valuable comments.
+The authors thank Julian Lange for useful literature input and Clemens Kreutz for valuable comments.
 The authors further thank Savanna Ratky for language corrections and Luzia Hanßum for support with formatting the manuscript.

### v3

@@
 To be reliable, they require careful planning of the design, execution, and reporting of the results.
-These issues have gained increasing attention in recent years, both in the context of simulation studies based on artificially generated data (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024) and in the context of real data-based benchmarking studies (Boulesteix et al., , 2017; Friedrich and Friede, , 2024).
+These issues have gained increasing attention in recent years, both in the context of simulation studies based on artificially generated data (Morris et al.,, 2019; Pawel et al., 2024b, ; Siepe et al.,, 2024) and in the context of real data-based benchmark studies (Boulesteix et al.,, 2017; Friedrich and Friede,, 2024).
 In this work, we focus on a particular aspect of the design of comparison studies that affects both simulation and benchmark studies, namely “method failure.” Method failure occurs when a method under investigation fails to produce an output for a (real or simulated) data set considered in the comparison study.
-The issue of method failure in comparison studies has begun to attract attention in recent years (Morris et al., , 2019; Pawel et al., , 2024; Siepe et al., , 2024), often focusing on non-convergence.
+The issue of method failure in comparison studies has begun to attract attention in recent years (Morris et al.,, 2019; Pawel et al., 2024b, ; Siepe et al.,, 2024), often focusing on non-convergence.
 However, there is little guidance on how to handle method failure appropriately, whether it occurs as non-convergence or in another form.
 In particular, published reviews have shown that method failure goes unreported in the majority of cases.
-For instance, in a systematic review of $42$ published simulation studies on the analysis of complex longitudinal patient-reported outcomes data, less than half “acknowledge that model non-convergence might occur(Hinds et al., , 2018).
+For instance, in a systematic review of $42$ published simulation studies on the analysis of complex longitudinal patient-reported outcomes data, less than half “acknowledge that model non-convergence might occur” (Hinds et al.,, 2018).
 Even more notably, only $12$ of $85$ applicable research articles reviewed by Morris et al., (2019) report convergence as a performance measure.
@@
 In predictive modeling, they are likely to compare the methods’ predictive performances (e.g., accuracy), that is, their ability to make predictions on a new set of observations.
-Here, the analysis is typically performed on real benchmarking data sets.
+Here, the analysis is typically performed on real benchmark data sets.
 For simplicity, we refer to both settings as “simulation studies” and “benchmark studies,” respectively (though admitting that there may be intermediate situations, e.g., simulation studies investigating predictive performance on artificial data).
@@
 4 0.78 0.76 NA
-Average accuracy ? 0.84 ?
+Av. accuracy ? 0.84 ?
 Regardless of the form in which a researcher encounters method failure, it leads to NA for the associated method-data set combination where otherwise an estimate (for simulation studies) or a performance value (for benchmark studies) is expected.
-Table 1 depicts a corresponding summary of the results of the hypothetical researcher’s study across all simulation repetition/benchmark data sets and methods.
+Table 1 depicts a corresponding summary of the results of the hypothetical researcher’s study across all simulation repetitions/benchmark data sets and methods.
 In simulation studies, the NAs complicate the derivation of performance (e.g., bias as the mean deviation of the estimates from the true value across all repetitions).
@@
 The rest of this paper is structured as follows.
-In Section 2, we give an overview of the status quo by reviewing and summarizing the different manifestations and handlings of method failure reported in published comparison studies, followed by a detailed discussion of why discarding data or imputing values are usually inadequate when encountering method failure.
+In Section 2, we give an overview of the status quo by reviewing and summarizing the different manifestations and handlings of method failure reported in published comparison studies, followed by a detailed discussion of why discarding data or imputing values is usually inadequate when encountering method failure.
 Building on this, Section 3 is a compilation of alternative recommendations for handling method failure, which differentiates between different scopes of the study and research stages in which it may take place.
-In particular, we promote using fallback strategies when encountering method failure, which enables the aggregation of performance across data sets even when method failure occurs and on top of that, reflects the behavior of real method users across many settings.
+In particular, we promote using fallback strategies when encountering method failure, which enables the aggregation of performance across data sets even when method failure occurs, and on top of that, reflects the behavior of real method users across many settings.
 Considering possible time constraints, we also discuss “slimmed-down” ways in which authors can implement part of our recommendations if unable to follow them completely.
@@
 Memory issues
-Remove method Modify implementation Change software Discard data for failing methods Discard data for all methods Impute Performance measure
+Remove method Modify implementation Change software Discard data for failing methods Discard data for all methods Impute Include performance measure
 van Smeden et al., (2016)a Simulation study to investigate why different published simulation studies offer conflicting minimal “events per variable (EPV)” recommendations for logistic regression.
@@
 Herrmann et al., (2021) Benchmark study on the predictive performance of thirteen survival time prediction methods based on eighteen high-dimensional multi-omics data.
-* $\times$ *     $\times$ $\times$
-Bischl et al., (2013) Comparison study of eleven classification methods on 49 real benchmark and simulated data sets.     * * *     $\times$ $\times$
-Nießl et al., (2022)b Meta-study to investigate the effect of different imputation procedures on the performance assessment and ranking of the methods.
+* $\times$ *     $\times$ b $\times$ $\times$
+Bischl et al., (2013) Comparison study of eleven classification methods on 49 real benchmark and simulated data sets.     * * *     $\times$
+Nießl et al., (2022)c Meta-study to investigate the effect of different imputation procedures on the performance assessment and ranking of the methods.
 Considered alternative procedures include imputation by the worst possible value and by a value that weighs the method’s overall proportion of method failure against its performance in the remaining folds.
@@
   • a
-    van Smeden et al van Smeden et al., (2016) observe that separated data sets (i.e., data sets leading to non-convergence) are commonly omitted from analyses.
+    van Smeden et al., (2016) observe in published comparison studies that separated data sets (i.e., data sets leading to non-convergence) are commonly omitted from analyses.
   • b
-    Nießl et alNießl et al., (2022) perform an example benchmark study based on the results of the study by Herrmann et alHerrmann et al., (2021) to illustrate how different ways of imputation can influence the results of a benchmark study.
+    Herrmann et al., (2021) apply a moderate form of data removal by restricting the analysis to the subset of smaller benchmark data sets for one method
+  • c
+    Nießl et al., (2022) perform an example benchmark study based on the results of the study by Herrmann et al., (2021) to illustrate how different ways of imputation can influence the results of a benchmark study.
     The authors did not report details on the manifestation of method failure.
@@
 One returns an error, while the other outputs the current, non-converged estimate in these instances (Zapf et al.,, 2024).
-Masaoud and Stryhn, (2010) encounter instances of “non-convergence or non-sensible estimates” (Masaoud and Stryhn, , 2010) for some method-data set combinations.
-In the study by Crowther et al., (2014), the methods are based on numerical integration and encounter estimation difficulties (non-convergence) with “challenging data sets(Crowther et al., , 2014).
-Hornung et al., (2024) report instances of method failure when methods are “not applicable” (Hornung et al., , 2024) to certain differences between training and test sets regarding the patterns of block-wise missingness.
+Masaoud and Stryhn, (2010) encounter instances of “non-convergence or non-sensible estimates” for some method-data set combinations.
+In the study by Crowther et al., (2014) the methods are based on numerical integration and encounter estimation difficulties (non-convergence) with “challenging data sets” (Crowther et al.,, 2014).
+Hornung et al., (2024) report instances of method failure when methods are “not applicable” to certain differences between training and test sets regarding the patterns of block-wise missingness.
 Additionally, they report calculation issues for one method explicitly disallowing certain structures in the training data set, resulting in failure.
-Calculation issues are reported by Ruxton and Neuhäuser, (2013), where some odds ratio estimation methods fail due to division by 00 when the $2\times 2$ contingency table contains one or more zero-entries.
+Calculation issues are also reported in the study by Ruxton and Neuhäuser, (2013), where some odds ratio estimation methods fail due to division by 0 when the $2\times 2$ contingency table contains one or more zero entries.
+Fernández-Delgado et al., (2014) report several data set characteristics associated with issues in required calculations, including “collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al.,, 2014).
+Additionally, a group of methods in their study requires a certain minimum number of observations per outcome class and is therefore not applicable to particularly small data sets.
 Another case of calculation issues can be found in the simulation study conducted by Dunias et al., (2024).
@@
 Some studies set fixed time budgets, so exceeding them is considered method failure.
-Others without fixed budgets encounter method failure when experiments for single methods run for an excessive amount of time (e.g.
-days), forcing the researcher to abort.
+Others without fixed budgets encounter method failure when experiments for single methods run for an excessive amount of time (e.g., days), forcing the researcher to abort.
 Unlike non-convergence and memory issues, method failure is actively declared by the researchers when runtime issues occur, as without these (naturally necessary) limitations, a method might eventually generate an output at some point.
-Resulting implications are explored in Section 3.1.6.
+Resulting implications are discussed in Section 3.1.6.
 Examples While Gijsbers et al., (2024) encounter the excess of predefined time budgets for model training, Herrmann et al., (2021) face runtime issues without fixed budgets.
@@
 Based on the published comparison studies, the handling of method failure can be commonly divided into two general approaches.
-Note that a literature review by Pawel et al., (2024), yielding a structured investigation of the frequencies of handling approaches from published simulation studies, results in four main handlings, which are among those we have commonly observed.
+Note that a literature review by Pawel et al., 2024a , yielding a structured investigation of the frequencies of handling approaches from published simulation studies, results in four main handlings, which are among those we have commonly observed.
 Resolving method failure Some authors resolve method failure completely so that no instances remain.
@@
 The first version of this approach, consisting of discarding data sets for the failing methods only, is followed by Zapf et al., (2024) for the method affected by errors due to non-convergence.
-Fernández-Delgado et al., (2014), comparing the methods’ performance based on two performance measures, discard the associated data sets for the methods affected by method failure in the comparison based on average accuracy.
+Fernández-Delgado et al., (2014), comparing the methods’ performances based on two performance measures, discard the associated data sets for the methods affected by method failure in the comparison based on average accuracy.
 The second version of discarding data, namely discarding the corresponding data sets for all methods, is applied by Hornung et al., (2024).
@@
 Above $20\%$, on the other hand, they impute by the worst possible value with the reasoning that the method’s behavior is “too unreliable for the current data set” (Bischl et al.,, 2013).
-A similar imputation strategy is employed by Herrmann et al., (2021).
+A similar imputation strategy is employed Herrmann et al., (2021).
 While using the same imputation rule when a method fails on more than $20\%$ of data sets, they otherwise impute by the mean performance of the method in the remaining “successful” data sets.
@@
 Instead, failure is the result we get!
-For this reason, we should be cautious about applying missing data approaches and particularly imputation to method failure.
+For this reason, we should be cautious about applying missing data approaches, and particularly imputation, to method failure.
 Note that this also includes measures of “weighted performance,” which weight the performance of a method according to the proportion of data sets with failure for that method.
@@
 However, we also highlight key considerations that can be addressed during the planning stage to help reduce both the potential for and unpredictability of failure in the first place.
+An overview of the recommendations can be found in Table 4.
 Using the terminology of Heinze et al., (2024), we distinguish between comparison studies in “early” and “late” research phases.
@@
 3.1 Practical recommendations for the handling of method failure
+Table 4: Checklist of recommendations for handling method failure.
+Recommended error handling workflow Section
+1. Investigate and understand the interplay underlying each instance of method failure
+Interplay may consist of data set characteristics, the method’s implementation, the given hardware, and software.
+3.1.1
+2. Check for the correct use of the method
+Distinguish between “use beyond original scope” and “method misuse” when detecting that a failing method is used differently than intended by its developers.
+3.1.2
+3. Check if data matches the scope of the study
+Check if the characteristics of the data associated with failure (and the failure itself) are realistic for the targeted users.
+3.1.3
+4. Be realistic when modifying method implementations
+For early-stage research. Restrict yourself to parameter adjustments that are realistic for future targeted users.
+3.1.4
+5. Consider a fallback strategy
+For late-stage research.
+Evaluate the performance of method pipelines of the type “use method A when it successfully produces an output, use method B otherwise.” Define the pipelines based on preliminary considerations and avoid selectively reporting the “best” pipelines.
+3.1.5
+6. Special case: runtime issues
+When real users running the affected method only on a single data set can be assumed to successfully obtain a model output, run the method on a preselected random subset of all data sets (selection best made in the planning stage).
+3.1.6
+7. Report failure proportions, underlying interplays, applied approaches, and share code
+Provide valuable insights to real users into the methods’ usability and enhance a deeper understanding of method failure in general.
+3.1.7
+8. Think ahead and pre-register, but remain flexible
+Conduct a pilot study before pre-registering the handling of method failure.
+Be aware that unforeseen failures may still occur; therefore, include error-handling mechanisms in your code from the start.
+3.1.8
+Slimmed-down error handling workflow when time constraints prevent recommendations 1-8
+Report (i) performance results after discarding the data sets for all methods, (ii) performance results after discarding the data sets for failing methods only, and (iii) failure proportions for all methods.
+Note that this approach does not provide a comparison of unconditional performance.
+3.2
 3.1.1 Consider method failure as the result of a complex interplay
@@
 Besides influencing runtime and memory consumption (Zimmermann,, 2020), data set characteristics contributing to method failure are highlighted by the observation that a method usually fails for specific data sets, not all of them.
-Certain data characteristics can complicate or completely prevent required calculations of a method within the maximum number of iterations (non-convergence).
-Recall the study of van Smeden et al., (2016) addressing perfect separation of the outcome variable by one or more predictor variables as a cause for non-convergence of logistic regression algorithms.
-Additionally, Fernández-Delgado et al., (2014), in their extensive comparison study of supervised classification methods, list “collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al., , 2014) as data set characteristics associated with method failure.
+Certain data characteristics can complicate or completely prevent required calculations of a method within the maximum number of iterations (non-convergence), time, or memory budgets.
+Note that this association is addressed in several comparison studies listed in Section 2.1.1.
+For instance, the study by van Smeden et al., (2016) addresses perfect separation of the outcome variable by one or more predictor variables as a cause of non-convergence of logistic regression algorithms.
+Recall also the data set characteristics associated with failure that are reported by Fernández-Delgado et al., (2014), including“collinearity of data, singular covariance matrices, and equal [predictor values] for all the training [observations], […] discrete predictor variables, classes with low populations, or too few classes” (Fernández-Delgado et al.,, 2014).
 Conversely, the association of method failure with the (implementation of) a given method becomes clear from the usual observation that not all methods fail for the same data sets.
@@
 In addition to the interplay of method and data set characteristics, however, software can also affect the (non-)failure of a method such that results are obtained in one software but not another.
-Consider, for instance, that Herrmann et al., (2021) reported that one of the methods under investigation was completely unusable due to an error in R that occurred using Windows as an operating system, while it did not occur under Ubuntu.
+Consider, for instance, that Herrmann et al., (2021) report that one of the methods under investigation is completely unusable due to an error in R that occurred using Windows as an operating system, while it does not occur under Ubuntu.
 The dependency on hardware, on the other hand, becomes clear in how working memory, besides being directly associated with possible memory issues, can also considerably affect runtimes.
@@
 Similarly, “too extreme” data may also affect real data-based studies.
-If one of the (real) benchmarking data sets is considered to be irrelevant to the scope of the study after careful examination, authors may consider removing the data set from the study completely.
+If one of the (real) benchmark data sets is considered to be irrelevant to the scope of the study after careful examination, authors may consider removing the data set from the study completely.
 See Section 4.2 for an illustration.
@@
 To be able to assess whether data structures associated with failure are too extreme, certain considerations should already be made in the planning stage.
-They should be aware of the scope and goal of their study: is it the goal to examine the general behavior of the method(s) (i.e., earlier stages of research) or to provide concrete recommendations for real-world practitioners (later stages of research)?
+Researchers should be aware of the scope and goal of their study: is it the goal to examine the general behavior of the method(s) (i.e., earlier stages of research) or to provide concrete recommendations for real-world practitioners (later stages of research)?
 Is the focus on testing the methods under regular or more extreme circumstances?
@@
 3.1.4 Be realistic when modifying method implementations
-In earlier stages of research, it may make sense to modify a method’s implementation entirely (i.e., for all simulation repetitions/benchmark data sets) when encountering method failure in the execution stage.
+In earlier stages of research, it may make sense to modify a method’s implementation entirely (i.e., for all simulation repetitions/benchmark data sets) when encountering method failure during execution.
 For instance, authors may conclude after execution that the default number of maximum likelihood iterations is generally too low, or that the given optimization algorithm may fail frequently, and that an alternative optimizer is more reliable.
@@
 On the other hand, it can sometimes be safely said that real (targeted) users, running the method on a single data set only, are likely to obtain an output.
-In these cases, when method failure leads to missing rather than undefined performance values, it may make sense to run the experiment on a randomly selected subset of all simulation repetitions/benchmarking data sets for the method affected by runtime issues (“failure by design”).
+In these cases, when method failure leads to missing rather than undefined performance values, it may make sense to run the experiment on a randomly selected subset of all simulation repetitions/benchmark data sets for the method affected by runtime issues (“failure by design”).
 To ensure that this “failure” is independent of data characteristics, it is important to select the subset of data independently of the data sets for which the method has particularly long runtimes.
-Researchers should therefore think about possible runtime problems and carry out the corresponding random selection of data sets in already the planning stage.
+Researchers should therefore think about possible runtime problems and carry out the corresponding random selection of data sets already in the planning stage.
 3.1.7 Report failure proportions, underlying interplays, applied approaches, and share code
@@
 This can serve as an important criterion for real-world users selecting a suitable method.
-Even the “best performing” method is of limited use if the amount of settings in which it can be applied without frequent failures is limited.
+Even the “best performing” method is of limited use if the number of settings in which it can be applied without frequent failures is limited.
 Ideally, the proportion of method failure should be included as an additional performance measure for all considered methods (see Morris et al., (2019)); see Zapf et al., (2024) for an example.
@@
 Adding to the previous paragraph, a corresponding reporting for each method can look like “method A, with a failure proportion of $X\%$, fails when a predictor variable is constant across all observations and when a binary predictor has highly imbalanced classes.” In addition to reporting the corresponding failure proportions, readers can learn in which settings a method is applicable, and no less importantly, in which it is not.
-Additionally, these insights serve as justification of the chosen handlings to method failure.
+Additionally, these insights serve as justification for the chosen handlings of method failure.
 An example of a corresponding reporting can be found in Hornung et al., (2024), who provide detailed descriptions in the supplement of their paper.
-Authors of comparison studies should additionally report their applied handlings.
-Any actions resulting from the above-described in-depth investigations and recommendations regarding method failure should be reported with sufficient detail.
-This includes removing any misused methods, removing data (or modification of DGMs) that do not match the scope of the study, any changes in implementations, and incorporating fallback strategies.
-Additionally, authors should justify their choices given the scope of the study (i.e., which users it is aimed at), and the research stage in which it takes place.
+Authors of comparison studies should additionally report their applied handlings with sufficient detail, that is, any actions resulting from the above-described in-depth investigations and recommendations regarding method failure.
+This includes removing any misused methods, removing data (or modifying DGMs) that do not match the scope of the study, any changes in implementations, and incorporating fallback strategies.
+Additionally, authors should justify their choices given the scope of the study (i.e., which users it is aimed at) and the research stage in which it takes place.
 The complexity of the occurrence and handling of method failure, which becomes clear from the previous sections, emphasizes all the more the value of sharing code, data, and intermediate results.
-Allowing authors of other, “new” comparison studies in-depth investigations to the specific instances of method failure in the given study may help them to a greater understanding of this topic–and enhance fruitful discussions on suitable handlings.
+Allowing authors of other, “new” comparison studies in-depth investigations of the specific instances of method failure in the given study may help them gain a greater understanding of this topic–and enhance fruitful discussions on suitable handlings.
 3.1.8 Finally: think ahead and pre-register, but remain flexible
@@
 In the following, we illustrate the guidelines presented in Section 3 through two fictive comparison studies featuring method failure: a statistical simulation study and a benchmark study at the interface between statistics and predictive modeling.
-Note that our focus is exclusively on the handling of undefined performance values and its consequences on the method comparison rather than the method comparison itself.
+Note that our focus is exclusively on the handling of method failure and its consequences on the method comparison rather than the method comparison itself.
 This means that both illustrations represent only a small excerpt (e.g., a few or a single scenario) of what would be a typical publishable study, exclusively focusing on those aspects of the studies that are relevant to the issue of method failure.
@@
 4.1.1 General
-This fictive statistical simulation study deals with the comparison of odds ratio estimation methods.
+This fictive statistical simulation study deals with the comparison of odds ratio (OR) estimation methods.
 Outcome and exposure are binary ($X,Y\in\left\{0,1\right\}$) and the observed number of subjects with $X=i$ and $Y=j$ is denoted as $n_{ij}$.
-The scenarios are defined based on different sample sizes $n.obs$, underlying true OR values, and probabilities of exposure $p_{x}$.
-Especially when the sample size is low, the underlying true OR is far from $1$, or $p_{x}$ is close to 00 or $1$, some simulated data sets may not contain any subjects for one or more of the four outcome-exposure combinations ($n_{ij}=0$).
-These “sampling zeros” cause calculation issues (e.g., division by 00) for many OR estimation methods, leading to either returning an error or, depending on the position of the sampling zero in the contingency table, non-meaningful OR estimations of 00 or $\infty$” for the corresponding data sets.
+The simulation scenarios are defined by the sample size $n.obs$, the underlying true OR, and the probability of exposure $p_{x}$.
+Especially when the sample size is low, the underlying true OR is far from $1$, or $p_{x}$ is close to 0 or $1$, some simulated data sets may not contain any subjects for one or more of the four outcome-exposure combinations ($n_{ij}=0$).
+These “sampling zeros” cause calculation issues (e.g., division by 0) for many OR estimation methods, leading to either returning an error or, depending on the position of the sampling zero in the contingency table, non-meaningful OR estimations of 0 or $\infty$” for the corresponding data sets.
 In practice, users often apply the “Haldane-Anscombe correction” (Haldane,, 1956; Anscombe,, 1956) to handle sampling zeros in the OR estimation.
-This correction consists of adding an offset of $+0.5$ to all four cells in the contingency table to avoid issues in computation.
+This correction consists of adding an offset of +0.5+0.5 to all four cells in the contingency table to avoid issues in computation.
 4.1.2 Design of the fictive study
-We vary the underlying true OR (OR $\in\{2,3,4,5\}$) and exposure probability $p_{x}$ ($p_{x}\in\{0.25,0.5\})$ while keeping the sample size fixed at $n.obs=50$.
-This results in eight distinct simulation scenarios (see Table 4 for an overview).
-For each scenario, we simulate 100 000100000100\,000100 000 outcome-exposure data sets.
+We vary the underlying true OR (OR{2,5}OR\ $\in\{$2,5\}) and exposure probability $p_{x}$ ($p_{x}\in\{0.25,0.5\})$ while keeping the sample size fixed at the relatively low value of $n.obs=50$.
+This results in four distinct simulation scenarios (see Table 5 for an overview).
+For each scenario, we simulate 100 000100\,000 outcome-exposure data sets.
 Five OR estimation methods are applied to the simulated data to obtain a point estimate for the true OR.
@@
 First, we compare the performance of the OR estimation methods when discarding data sets with sampling zeros for all methods versus only for the failing methods (“ad hoc approaches”; Section 4.1.3).
-Though users often apply the Haldane-Anscombe correction when encountering sampling zeros, this first part illustrates how the preference of some methods may mainly be driven by the way data sets with sampling zeros are discarded.
+Though real users often apply the Haldane-Anscombe correction when encountering sampling zeros, this first part illustrates how the preference of some methods may mainly be driven by the way data sets with sampling zeros are discarded.
 Second, we repeat the analysis implementing the principle of fallback strategies suggested in Section 3.1.5.
@@
 For instance, for the manual estimation of the OR (method “Manual”), only the Haldane-Anscombe correction is applied as a fallback, assuming that a user applying the manual computation in the first place, which is very easy to implement, is unlikely to resort to a method that requires notably more effort.
-On the other hand, for methods Midp and Fisher for data sets with sampling zeros, we (i) either apply the Haldane-Anscombe correction, (ii) use method Small, or (iii) use method Woolf, assuming that corresponding users are more likely to employ fallbacks that are slightly more difficult to apply.
+On the other hand, for methods Midp and Fisher for data sets with sampling zeros, we either (i) apply the Haldane-Anscombe correction, (ii) use method Small, or (iii) use method Woolf, assuming that corresponding users may also employ fallbacks that are slightly more difficult to apply.
 Together with Woolf and Small, this leads to nine “methods” (“pure” methods or pipelines) being compared in total.
@@
 4.1.3 Occurrence of method failure and ad hoc handlings
-Table 4 displays the proportion of the 100 000100000100\,000100 000 simulated data sets containing at least one sampling zero across the eight simulation scenarios.
-As expected, this proportion increases with an increasing true OR$>1$, while decreasing when exposure $X$ is more balanced.
-Table 4: Overview of the proportion of the 100 000100000100\,000100 000 simulated data sets with sampling zeros across the simulation scenarios.
-$n.obs$: number of observations, $p_{x}$: exposure probability.
-Simulation scenario n.obs true OR $p_{x}$ Sampling zero proportion
-1 50 2 $0.25$ 1.28%
-2 50 3 $0.25$ 3.96%
-3 50 4 $0.25$ 7.8%
-4 50 5 $0.25$ 12.0%
-5 50 2 $0.5$ 0.01%
-6 50 3 $0.5$ 0.11%
-7 50 4 $0.5$ 0.50%
-8 50 5 $0.5$ 1.27%
+Table 5 displays the proportion of the 100 000100\,000 simulated data sets containing at least one sampling zero across the four simulation scenarios.
+As expected, this proportion increases with an increasing true OR>1>1, while decreasing when exposure $X$ is more balanced.
+Table 5: Overview of the proportion of the 100 000100\,000 simulated data sets (fixed sample size $n.obs=50$) with sampling zeros across the simulation scenarios. $p_{x}$: exposure probability.
+Simulation scenario true OR $p_{x}$ Sampling zero proportion
+1 2 $0.25$ 1.28%
+2 5 $0.25$ 12.0%
+3 2 $0.5$ 0.01%
+4 5 $0.5$ 1.27%
 Figure 1: Performance ranks of the OR estimation methods across the eight simulation scenarios based on empirical bias on the $\log$-scale, where a rank of 1 means that the method has the lowest empirical bias in the given scenario.
@@
 (B) The ranks are compared after applying fallback strategies to methods Midp, Fisher, and Manual.
-For Midp and Small, the fallbacks Small, Woolf, and the Haldane-Anscombe correction (“$+0.5$”) are used.
+For Midp and Small, the fallbacks Small, Woolf, and the Haldane-Anscombe correction (“+0.5+0.5”) are used.
 For Manual, only the Haldane-Anscombe correction is used as fallback.
@@
 Abbreviation of methods: “Fis”: Fisher; “Mid”: Midp“; “Sma”: Small; “Woo”: Woolf; “Man”: Manual (i.e., manual estimation of the OR).
-(B): The combination of original method $m_{1}$ and fallback $m_{2}$ is marked by “$m_{1}/m_{2}$.”
-Ad hoc approaches  For the first analysis, Figure 1(A) shows that the way of discarding data sets with sampling zeros can substantially influence the results of the method comparison.
-This can lead to changes in the performance rank by up to three positions in some scenarios (see method Woolf) and particularly affect which method is chosen as the best.
+(B): The combination of original method $m_{1}$ and fallback $m_{2}$ is marked by $m_{1}/\penalty 50m_{2}$.”
+Ad hoc approaches In the first analysis, Figure 1(A) demonstrates that the way of discarding data sets with sampling zeros can substantially affect the results of the method comparison, with discrepancies regarding performance ranks already observable at a sampling zero proportion of just over $1\%$.
+This impact becomes more pronounced as the proportion of sampling zeros increases, making a difference in the observed performance by up to three ranks.
+In particular, it can affect which method is perceived as best-performing.
 Note that a more detailed investigation of Woolf’s performance shows that for the data sets without sampling zeros, Woolf slightly underestimates the true OR, while for the few data sets with sampling zeros, Woolf notably overestimates it.
 This leads to top performance when all data sets are included in its performance assessment (while discarding data sets with sampling zeros for the failing methods).
-Generally, the manner of discarding data sets has a larger impact on the performance comparison for scenarios with a higher proportion of sampling zeros.
-However, differences in rankings within a given scenario are already observable with a proportion of $0.5\%$ of data sets with sampling zeros.
 4.1.4 Recommended handling: using fallbacks
@@
 Since this applies to both fallbacks, the Haldane-Anscombe correction and Small, it is debatable which fallback is preferable.
-Both fallbacks for Midp frequently switch ranking positions across the performance measures and simulation scenarios, however, fallback Small might be even easier to implement than the Haldane-Anscombe correction, requiring only a single parameter modification from Midp.
+Both fallbacks for Midp frequently switch ranking positions across the performance measures and simulation scenarios, however, fallback Small might be even easier to implement than the Haldane-Anscombe correction, requiring only a simple code modification from Midp.
 4.2 Example 2: Comparing methods for computing confidence intervals for the generalization error
@@
 The fictive study aims to compare the performance of two methods for constructing confidence intervals (CIs) for the generalization performance of a prediction model based on resampling-based estimates, such as cross-validation.
-The “naive” method “$N$” for constructing the CIs ignores the dependence between the generalization performances in the resampling repetitions, while method “$C$” corrects for it.
+The “naive” method “NN” for constructing the CIs ignores the dependence between the generalization performances in the resampling repetitions, while method “CC” corrects for it.
 The performance of both methods is assessed in terms of coverage.
@@
 interval
-Method $N$ $0.54$ $0.54$ $0.38$ $0.68$
-Method $C$ $1$ $1$ $1$ $1$
+Method NN $0.54$ $0.54$ $0.38$ $0.68$
+Method $C$ 11 11 11 11
 The coverages resulting from the ad hoc approaches are displayed in Table 6 (a).
@@
 This illustrates well that the further the investigations are taken, the apparent adequate handling of method failure can change notably.
-In this case, based on the information gained, the adequate handling depends on the aim of the study and requires further considerations such as the following.
+In this case, based on the information gained, adequate handling depends on the aim of the study and requires further considerations such as the following.
 A prediction problem where splitting often does not occur can be viewed as an edge case.
 If the study’s scope includes such edge cases, one could attempt to suitably modify the parameters of the classification tree to favor splitting.
-If, however, the study aims to compare both methods in regular scenarios, one might consider adding additional predictors to favor splitting more strongly, resort to an alternative prediction model (e.g., logistic regression), an alternative measure of generalization performance (e.g., accuracy), or use a different benchmarking data set.
+If, however, the study aims to compare both methods in regular scenarios, one might consider adding additional predictors to favor splitting more strongly, resort to an alternative prediction model (e.g., logistic regression), an alternative measure of generalization performance (e.g., accuracy), or use a different benchmark data set.
 5 Discussion
@@
 This means imputation, a common but missing-data-inspired approach to handling method failure, is usually inadequate.
-The same applies to the popular and seemingly logical approach of discarding data associated with failure (either for all or failing methods), which almost never enables the comparison of unconditional performance.
+The same applies to the popular and seemingly logical approach of discarding data associated with failure (either for all or the failing methods), which almost never enables the comparison of unconditional performance.
 In this work, we provide alternative approaches that might be more appropriate to most occurrences of method failure.
-An approach that in our view deserves more attention than it receives currently in the literature consists of evaluating whole “pipelines,” including one or several fallback strategies in case of method failure.
+An approach that in our view deserves more attention than it receives currently in the literature consists of evaluating whole “pipelines” of methods, including one or several fallback strategies in case of method failure.
 This approach reflects the typical behavior of most users of methods in practice while enabling the comparison of unconditional performance if constructed carefully: if one method does not work, they resort to an alternative one.
@@
 They lead to helpful strategies for dealing with it, as outlined as part of our recommendations.
-However, we also believe that method failure is only one of many issues related to comparison studies that deserve much more time, efforts, and methodological strength towards more reliable evidence on the behavior of methods in the context of a “replication crisis in methodological research” (Boulesteix et al., , 2020).
+However, we also believe that method failure is only one of many issues related to comparison studies that deserve much more time, effort, and methodological strength towards more reliable evidence on the behavior of methods in the context of a “replication crisis in methodological research” (Boulesteix et al.,, 2020).
 We need more well-designed neutral comparison studies (Boulesteix et al.,, 2017)—late phase studies according to the terminology of Heinze et al., (2024).